A Complete Bibliography of Publications in the Intel Technology Journal

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

27 September 2023 Version 1.13

Title word cross-reference

.NET [135].

/Fortran [99].

10 [179, 178]. 130nm [104].

2.0Volts [10]. 21st [9, 24].

300mm [105, 21]. 30nm [108]. 3945ABG [268]. 3D [41].

4 [246, 72, 73, 168, 171, 75, 76, 170, 169, 77].

60nm [104]. 64 [26, 46, 45, 27, 28, 38, 44]. 64-bit [48].

802.11b [49]. 802.16 [190, 189].

90nm [168, 171, 170]. 915GMS [210]. 945GMS [267].

AAL2 [119]. Abstractions [137]. Access [190]. Achieving [115, 19]. Across [84, 204, 6]. Adaptive [199]. Adding [116]. Addressing [66, 160]. Advanced [162, 16, 241, 247, 254]. AdvancedTCA [159, 158]. AdvancedTCA* [254]. AdvancedTCA/CGL [159]. Advancements [202]. Advantage [159]. Air [153]. Algorithms [14]. All-IP [118]. Always [55]. Analysis [223, 220, 12, 16, 173, 114, 247, 170]. Antenna [146, 191]. AOAC [55]. Applicability [197]. Application [204, 179, 60, 158]. Applications [42, 217, 100, 25, 202, 198, 180, 117, 277, 135]. Architectural [62]. Architecture [48, 64, 204, 26, 172, 80, 254, 161, 65, 27, 23, 263, 28, 269, 56, 231, 96, 3, 34, 206, 83, 268, 38, 44, 271, 125, 29, 41]. Architectures [158]. Area [154]. Arithmetic [46]. Array [246, 61]. ASP [135]. Assembly [54, 38]. Assessments [58]. Asset [51]. Atlantic [84]. Audio [88, 213, 35]. Automated [179]. Automation [236, 87]. Autonomic [198].

Balance [77]. Band [179]. Bands [86]. Based [73, 223, 221, 163, 172, 254, 23, 183, 58, 264, 243, 43, 256, 267, 77, 68]. Bayes [205]. Beginning [261]. Benchmarking [279]. Benefit [173]. Best [134, 259]. Best-in-Class [259]. Better [210, 230]. Beyond [204, 108]. Binary [46]. Bioinformatics [222]. bit [48]. Blocks [165, 118]. Bluetooth [62, 55, 56, 57]. Bringing [200]. Brings [142]. Broadband [187, 80, 185]. Budgeting [145]. Build [257]. Building [194, 165, 118]. Built [216, 266, 215]. Business [65, 69, 206, 51].

C [99, 179]. C-Band [179]. CAD [47, 36]. Campus [255]. Capabilities [9]. Capacity [233, 51, 86, 230]. Case [220, 185, 255]. Cell [172]. Cell-Based [172]. Cellular [50]. Center [164, 65]. Centers [66, 52]. Centrino [266, 261, 267, 259]. Centrino^{T M} [140, 207, 216, 141, 213, 212, 215, 209, 142]. Century [9, 24]. Certifying [52]. CGL [159]. Chain [234, 67, 232, 231]. Challenges [246, 120, 188, 47, 172, 106, 244, 239, 160, 224, 59, 24, 70]. Chaos [53]. Characteristics [158]. Characterization [279, 11]. Chip [54, 181, 61, 78]. Chip-Join [54]. Chipset [210, 267]. Circuit [188, 30]. Class [259]. Classes [258]. Client [201, 276]. CMOS [192, 107]. CMP [264]. Cockpit [67]. Coexistence [86]. Collaboration [259, 206, 256]. Collaboration-Enabling [259]. Comes [212]. Common [214]. Communication [32, 146, 269, 248, 83]. Communications [157, 155, 81, 253, 156, 277, 175, 86, 258]. Compiler [133, 45, 85, 99]. Compilers [169]. Compliance [37]. Components [214]. Computation [28]. Compute [217, 101]. Compute-Intensive [217, 101]. Computer [88, 22, 221, 220, 93]. Computers [151]. Computing [164, 194, 254, 55, 36, 261, 224]. Conferencing [35, 18]. Configuring [198].

Connected [55]. Connection [268]. Connectivity [64, 154]. Consumer [79]. Consumption [214, 74]. Content [53, 121, 127]. Control [161, 5, 190]. Converged [253, 252]. Convergence [164, 79, 250]. Conversion [183]. Cooling [243]. Copy [7]. Core [260, 263, 264, 265, 261]. Corporate [53]. Corporation [226]. Correctness [14]. Cost [179, 173]. Cost/Benefit [173]. CPE [80]. Creating [149]. Creator [90]. Cross [36]. Cross-Platform [36]. CSP [83]. Cu [104]. Current [103]. Cycle [6].

Data [164, 223, 249, 222, 65, 202, 66, 252, 52]. Data-Mining [223, 222]. Datapath [47]. Decision [67]. Decoder [2]. Defect [5, 43]. Defect-Based [43]. Delivery [242, 80]. Demand [226, 227, 152]. Demands [128]. Demystifying [18]. Deployment [150]. Design [47, 172, 91, 12, 30, 36, 76, 66, 71]. Designs [34]. Developing [89, 198, 135]. Development [47, 238, 239, 36, 105, 87, 44, 78]. Device [258]. Devices [151, 138]. Diagnostics [205]. Dielectrics [104]. Digital [2, 121, 90, 128, 107, 123, 127, 258]. Directed [274]. Directions [64, 110]. Discussion [77]. Dissection [93]. Distributed [161]. Distribution [121, 124]. Divide [14]. Dolby [2]. Duo [260, 263, 266, 264, 265, 261, 267]. During [59]. Dynamic [133, 152].

e-Business [65, 69, 51]. e-Home [79]. e-Procurement [232]. Early [78]. eBusiness [60]. Edge [164]. Efficient [73, 273]. Electrical [71]. Electronic [238, 239]. Electronics [79]. Elements [161, 108, 107]. Embedded [277]. Emergence [211]. Emergent [258]. Emerging [178, 110, 185, 225]. Enable [230]. Enabled [141]. Enabler [43]. Enabling [259, 55, 34]. Engineering [60]. Enhanced [68]. Enhancement [40]. Enterprise [164, 196, 204, 194, 278, 177, 201, 199, 200, 252, 225, 203, 134, 197]. Entity [236]. Environment [136, 132, 30, 36, 44, 40]. Environments [133, 138]. Equipment [231]. Era [175, 261]. Estimation [145]. Ethnographic [84]. ETOX^{T M} [106]. Europe [84]. EUV [8]. Evolution [50, 64, 65, 253, 63, 7]. Evolving [87]. EXACTLY [7]. Experience [259, 125]. Experiences [256]. Exploitation [73]. Exploring [98]. Express [267]. Express* [211]. Extending [187, 275, 82]. Extensions [42, 40, 31]. External [149].

Fabrics [158]. Face [22]. Factor [267]. Factory [67, 6, 230]. Failure [16, 247]. Family [162]. Fault [247]. Feature [25]. Featuring [104]. Finding [244]. Flash [106]. Flexible [132]. Flip [54, 61]. Flip-Chip [54, 61]. Floating [14, 46, 37]. Floating-Point [14, 46, 37]. Flows [36]. Forecasting [229]. Foreword [194, 140, 95, 112, 239, 227, 218, 156, 250, 175, 261, 131, 185, 148, 167]. Form [267]. Formally [37]. Fortran [99]. Forward [271]. Foundation [203]. Framework [53, 162, 198, 117, 206, 123, 29]. Freeing [125]. Frequencies [192]. Full [179, 173]. Fully [192]. Functions [28]. Fundamentals [178]. Future [246, 64, 103, 248, 247].

Game [91]. Gate [108]. Gb [179, 178]. Gb/s [179, 178]. Generation [113, 50, 207, 216, 213, 138, 257, 211, 212, 243, 215, 209]. Getting [69]. Gigabit [115]. Gigabit/sec [115]. Gigahertz [34]. Global [187, 206]. Globalization [232]. Goes [219]. Graphics [212, 142]. Grid [246, 61].

H.323 [11, 33, 18]. Hand [151]. Hand-Held [151]. Handheld [137]. Hardware [273, 37]. Held [151]. High [2, 242, 246, 153, 120, 132, 183, 128, 107, 212, 241, 105, 124]. High-Performance [242, 120, 132, 107, 212, 241, 105]. High-Quality [124]. High-Speed [183, 128]. High-Throughput [153]. High-Volume [246]. Highly [217]. History [92]. Hold [173]. Hold-Scan [173]. Home [79, 121, 126, 129, 128, 82, 123, 124, 127, 258, 122]. Hop [128]. Hotspot [150]. HW [268]. Hyper [97, 100, 95, 101, 96, 169, 99]. Hyper-Threading [97, 100, 95, 101, 96, 169, 99].

I/O [274, 125]. IA [48, 163, 26, 46, 45, 27, 28, 160, 38, 44]. IA- [163]. IA-64 [48, 26, 46, 45, 27, 28, 38, 44]. Idea [89]. IEEE [14, 46, 190, 37, 49, 189]. II [1]. III [73, 25, 39, 40, 41]. Illinois [23]. Illinois-Intel [23]. Immunity [76]. Impact [74, 101]. Implementation [2, 39, 264, 99]. Implementing [120, 69, 119]. Improving [6]. Increasing [74]. Indirect [232]. Indium [183]. Indoor [86]. Industry [69]. Information [203]. Infrastructure [34, 277, 122]. Initiation [253]. Initiative [69]. Innovation [149, 142, 149]. Innovations [112]. Input [213]. Input/Output [213]. Integrated [80, 237, 192, 34, 142]. Integrating [57]. Integration [106, 107, 180, 203]. Intel [274, 113, 233, 48, 72, 64, 221, 10, 194, 236, 226, 260, 275, 45, 259, 65, 155, 62, 19, 23, 263, 257, 15, 60, 266, 227, 20, 7, 264, 13, 265, 273, 277, 261, 276, 268, 17, 109, 255, 267, 77, 271, 78]. Intel(R) [88, 90, 95, 91, 92, 85, 93, 87, 99, 168, 210, 140, 207, 216, 141, 213, 171, 145, 143, 212, 215, 144, 170, 169, 209, 142, 165]. Intensive [217, 101]. Interconnect [244, 76, 180, 71]. Interconnects [181, 104]. Interface [153, 22, 215]. Internal [149]. International [226, 227]. Internet [64, 32, 201, 66, 31, 18]. Internetworking [129, 150]. Interoperability [254, 123]. Interoperable [187, 122]. Intranets [53]. Introduction [32, 263]. Inventory [235]. Investment [117]. IP [33, 66, 160, 256, 118]. IPv6 [129]. Isolation [247]. issue [4, 48]. Issues [66]. Itanium [85]. Iterative [14]. ITJ [130, 139, 95, 112]. IXA [117, 165]. IXP [113, 163, 162, 160]. IXP-Based [163]. IXP2800 [115].

Java [134]. Join [54]. Journal [155].

Kernels [26, 27]. Key [43].

Lab [92]. LAN [153, 150]. Land [246]. Land-Grid-Array [246]. Landing [60]. Language [38]. Languages [162]. Large [223, 19, 224]. Large-Scale [223, 224]. Latency [98]. Layer [189]. Layout [30]. Leadership [194]. Learning [221, 202]. Learning-Based [221]. Learning/Data [202]. Length [108]. Level [71, 243, 83, 144]. Library [221, 172, 23]. Life [6]. Life-Cycle [6]. Line [5]. Lithography [8, 109]. Local [154]. Logic [10, 104]. Looking [84]. Low [213, 179, 104, 142]. Low-Cost [179]. Low-K [104]. Low-Power [213]. LVS [171].

M [145, 143, 144]. Mainstream [219, 271]. Make [84]. Making [85]. Managed [133, 132, 138, 131]. Management [67, 51, 19, 60, 266, 201, 265, 215, 203, 68, 245]. Managing [226, 74, 229, 227, 68]. Manufacturing [103, 202, 12, 9, 105, 87]. Material [215]. Materials [245]. Maximize [5]. Me2Cam* [91]. Mechanism [58]. Mechanism-Based [58]. Mechanisms [35]. Media [100, 257, 124]. Medium [81, 190]. Medium-Range [81]. Meet [260]. Meeting [128]. Memory [106, 266]. Merging [249]. Metadata [203]. Method [7]. Methodologies [120]. Methodology [114, 58]. Methods [84, 40]. Metrics [163]. Micro [243]. Microarchitecture [168, 143, 75, 1, 96]. Micron [10, 17]. Microprocessor [74, 63, 87, 59]. Microprocessors [242, 16, 1, 173, 105]. Microscope [93]. Migration [43]. Millimeter [192]. MIMO [269]. Mining [223, 222, 202, 218]. Mixed [107]. Mixed-Signal [107]. MMX [2, 4, 1, 3]. Mobile [151, 210, 140, 214, 207, 216, 141, 213, 138, 259, 266, 57, 211, 212, 261, 215, 267, 209, 142]. Mobility [187]. Modeling [235]. Models [136, 223, 214, 216, 276]. Modular [157]. Morpher [88]. MOS [24]. Movie [90]. Multi [128, 261]. Multi-Core [261]. Multi-Hop [128]. Multicarrier [146]. Multimedia [11, 35, 199, 33, 256, 18]. Multiple [191]. Multiprocessor [23]. Multithreading [23, 98].

Nano [243]. Network [113, 205, 120, 163, 161, 126, 116, 112, 227, 33, 114, 198, 117, 268, 252, 119, 68, 118]. Network-One [252]. Networking [165]. Networks [187, 154, 35, 177, 183, 128, 199, 185, 152, 29]. Next [113, 216, 138, 257, 211, 243]. Next-Generation [216, 138, 257, 243]. Nike [34]. Noise [76]. Number [25].

O [274, 125]. OFDMA [189]. On-Chip [181]. One [252]. Open [221, 80, 231, 149, 132]. OpenMP [99]. Operating [26, 27, 19]. Operations [67, 46, 12]. Opportunities [239]. Optical [8, 179, 177, 183, 178, 181, 180, 175]. Optimization [233, 145, 224]. Optimizing [234, 135]. Optimum [71]. Options [230]. Optoelectronic [183]. Organic [61, 245]. Oriented [204]. Our [230]. Output [213, 212]. Overview [45, 62, 16, 56, 20, 3, 49].

Package [238, 244, 239, 248, 241, 58, 247, 71, 243]. Package-Level [71, 243]. Packages [61, 245]. Packaging [54, 179, 63, 110]. Packet [115, 164, 35]. Parallel [217, 224]. Parallelism [73, 19]. Pattern [90]. PC [79, 216, 141, 82, 256, 77]. PC-Based [256]. PCI [211]. Peak [41]. Pentium [72, 246, 73, 25, 75, 16, 1, 39, 76, 77, 40, 41]. Pentium(R) [168, 171, 145, 143, 144, 170, 169]. Perceptual [22]. Performance [242, 120, 223, 163, 279, 214, 222, 134, 132, 35, 143, 15, 107, 114, 212, 241, 105, 71, 170, 99, 70, 40, 41]. Personal [32]. Phoenix [29]. Phosphide [183]. Phosphide-Based [183]. Photonics [182]. Physical [108, 189]. Pin [61]. Plane [161]. PlanetLab [197]. Planning [233, 234, 51, 229, 30, 21]. Platform [157, 214, 216, 149, 213, 36, 212, 13, 225, 267, 77, 209, 142, 135, 125, 132]. Platforms [210, 141, 266, 211, 215, 271, 165]. Play [88, 90, 91, 92, 93]. Point [14, 46, 37]. Policy [68]. Policy-Based [68]. Portability [117]. Portal [53]. Porting [26, 27]. Power [242, 214, 213, 145, 74, 266, 265, 142]. Practical [29]. Practices [134]. Pragmatic [68]. Pre [97, 26, 27, 44]. Pre-Silicon [97, 26, 27, 44]. Precomputation [98]. Preface [102, 111, 94, 130, 139, 147, 166, 174, 184, 193, 217, 238, 226, 207, 270, 260, 249, 155]. Preserving [117]. Principles [91]. PRO [268, 16]. PRO/Wireless [268]. Proactive [196, 199, 197]. Proactively [200]. Problems [223, 224]. Process [10, 107, 105, 206, 17]. Processes [233]. Processing [115, 164, 163, 155, 156, 160]. Processor [246, 72, 73, 168, 260, 171, 85, 116, 25, 145, 143, 263, 75, 39, 76, 114, 58, 264, 117, 265, 273, 144, 170, 169, 119, 77, 118, 40, 41]. Processor-Based [73, 77]. Processors [113, 120, 112, 1, 160]. Procurement [232]. Product [89]. Products [57]. Programmable [29]. Programming [38, 40]. Proposed [258]. Protection [116, 127]. Protocol [253]. Proving [14]. Provisioning [190]. Public [150]. Purposes [26, 27].

Q1 [94, 130, 95]. Q2 [139]. Q3 [112]. Quality [2, 124, 17, 255]. Quarter [17]. QX3 [93].

Radios [192]. Ramps [230]. Range [81]. Ray [219]. Reach [180]. Reactive [35]. Real [237, 256]. Real-Time [256]. Recognition [218]. Recollections [78]. Recovering [35]. Redefining [279, 231]. Release [60]. Reliability [244, 58, 52, 17]. Remainder [14]. Remote [125]. Requirements [225]. Resource [19]. Resume [201]. Reviewers [176, 186, 195, 208, 228, 240, 272, 251, 262]. Revolution [231]. RF [188, 192]. RFIC [268]. RFID [237]. Roadmap [109]. Role [87]. Root [14]. RosettaNet [236, 69]. Runtime [133, 136, 132, 137, 138, 131].

s [179, 178]. Scalability [223, 222]. Scalable [199, 13, 83, 165, 189]. Scale [223, 19, 224]. Scaling [106, 24]. Scan [173]. Seamless [154, 259]. sec [115]. Second [207, 213, 212, 215, 209]. Second-Generation [207, 213, 212, 215, 209]. Security [136, 126, 116, 200, 49]. Selection [146, 215]. Self [198]. Self-Configuring [198]. Semiconductor [103, 202, 9]. Sense [84]. Serial [25]. Server [279]. Servers [257]. Service [204, 190, 52].

Service-Oriented [204]. Services [249, 80, 33, 198, 13, 83, 68]. Session [253]. Set [18]. Short [81, 180]. Short- [81]. Signal [107]. Silicon [164, 97, 182, 70, 26, 27, 44]. SIMD [31, 42, 40]. Simultaneous [54]. SIP [253]. Small [267]. Smaller [210]. Smart [205, 89, 92]. SoftSDV [44]. Software [88, 62, 162, 34, 117, 44, 41]. Solution [80]. Solutions [244, 34, 243, 224, 225, 77]. SONET [115]. Sound [88]. Source [221]. Space [137, 66]. Special [4, 48]. Speculative [98]. Speed [183, 128]. Square [14]. Stack [62, 41]. Standard [46]. Standards [50, 254, 18, 258]. Standards-Based [254]. StarJIT [133]. Statistical [202]. Storage [213]. Story [237]. Streaming [42, 31, 40]. Streams [11]. Strengthening [232]. Structural [43]. Study [220, 255]. Successful [204, 43]. Successor [8]. Supercomputer [19, 15, 20, 13]. Supplementary [33]. Supply [234, 67, 226, 232, 231, 227]. Supply-Chain [234, 232]. Support [67, 23, 273, 169]. Surveillance [220]. Suspend [201]. Suspend/Resume [201]. Switched [35]. Synthesis [218]. System [188, 26, 91, 27, 19, 12, 266, 180, 83, 268, 144]. System-Level [83, 144]. Systems [73, 220, 146, 23, 248, 173, 241, 264, 191, 70, 163].

Taking [271]. TCO [159]. TCP [160]. TCP/IP [160]. Team [206]. Technical [176, 186, 195, 208, 228, 240, 272, 251, 262, 64]. Technique [215]. Techniques [16, 247]. Technologies [147, 213, 257, 177, 248, 178, 110, 241, 247, 131, 148, 245]. Technology [2, 187, 4, 48, 246, 168, 103, 10, 97, 140, 238, 207, 100, 216, 141, 213, 90, 95, 171, 275, 54, 259, 106, 155, 179, 55, 81, 232, 129, 239, 266, 1, 57, 212, 61, 101, 96, 7, 3, 261, 243, 191, 215, 170, 169, 267, 209, 104, 99, 142, 274, 273, 277, 276, 271]. Technology-Based [243]. Telecom [254]. Telephony [33]. Terminal [11]. Test [231, 43]. Testing [59]. TFLOPS [19, 15, 20, 13]. Thermal [266, 265, 215, 59, 70]. Thermomechanical [245]. Third [50]. Third-Generation [50]. Threading [97, 100, 95, 101, 96, 169, 99]. Throughput [153, 6]. Time [256]. Tomorrow [65]. Tool [67]. Tools [162, 38, 98]. Toy [92]. Toys [89, 92]. Tracing [219]. Tracking [22]. Tradeoffs [39]. Trading [236]. Transceivers [179, 178]. Transcendental [28]. Transfer [7]. Transforming [53]. Transistor [108, 24]. Transistors [104]. Transition [21]. Transmitter [179]. Tree [223]. Tree-Based [223]. Trends [246]. Triplets [249]. Tunable [179]. Tuning [42]. TV [212].

Ultra [81, 180]. Ultra-Short-Reach [180]. Ultra-Wideband [81]. Uncertainty [229]. Underfill [54]. Understanding [225]. Union [151]. Unlicensed [86]. UPnP* [125]. Usage [214, 216, 276]. Use [22, 98]. User [22]. Uses [217, 258]. Using [84, 129, 257, 160, 276, 18, 230, 40, 2].

Validating [72]. Validation [97, 26, 27, 145, 144, 170]. Value [149]. Verifying [37]. via [116]. Video [249, 220, 252]. View [68]. Virtual [91]. Virtualization [274, 279, 275, 278, 273, 277, 276, 271]. Vision [22, 221, 220].

Voice [249, 252, 119]. VoIP [259, 255]. Volume [246]. VTune [40].

Wafers [105]. Wave [192]. Wavelength [183]. Web [135]. Wideband [81]. WiMAX [187, 188, 185, 191]. Wired [152]. Wireless [187, 153, 79, 147, 154, 137, 81, 269, 150, 248, 128, 199, 268, 252, 86, 148, 152, 118]. WirelessMAN [189]. WLAN [151, 268]. Workload [220]. Workloads [222, 101]. WPAN [151].

Xen [275].

Yesterday [65]. Yields [5].

References

Kagan:1997:MMP

[1] Michael Kagan, Simcha Gochman, Doron Orenstien, and Derrick Lin. MMX microarchitecture of Pentium processors with MMX technology and Pentium II microprocessors. Intel Technology Journal, (Q3):8, 1997.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/q31997/articles/art_4.htm;   http://developer.intel.com/
technology/itj/q31997/pdf/micro.pdf.

Abel:1997:IHQ

[2] James C. Abel and Michael A. Julier. Implementation of a high quality Dolby digital decoder using MMX technology. Intel Technology Journal, (Q3):11, 1997. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q31997/articles/art_3.htm;   http:/
/developer.intel.com/technology/itj/q31997/pdf/decoder.pdf.

Mittal:1997:MTA

[3] Millind Mittal, Alex Peleg, and Uri Weiser. MMX technology architecture overview. Intel Technology Journal, (Q3):12, 1997. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q31997/
articles/art_2.htm;   http://developer.intel.com/technology/
itj/q31997/pdf/archite.pdf.

Anonymous:1997:SIM

[4] Anonymous. Special issue on MMX technology. Intel Technology Journal, (Q3), 1997. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/q31997.htm.

Mittal:1998:LDC

[5] Sanjiv Mittal and Peter McNally. Line defect control to maximize yields. Intel Technology Journal, (Q4):3, 1998. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q41998/
articles/art_6.htm;   http://developer.intel.com/technology/
itj/q41998/pdf/defect_control.pdf.

Kempf:1998:ITA

[6] Karl G. Kempf. Improving throughput across the factory life-cycle. Intel Technology Journal, (Q4):6, 1998. ISSN 1535-766X. URL http://

developer.intel.com/technology/itj/q41998/articles/art_5.htm;
 http://developer.intel.com/technology/itj/q41998/pdf/factory_
life_cycle.pdf.

McDonald:1998:EIC

[7] Chris J. McDonald. The evolution of Intel’s Copy EXACTLY! technology transfer method. Intel Technology Journal, (Q4):6, 1998. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q41998/
articles/art_2.htm;   http://developer.intel.com/technology/
itj/q41998/pdf/copyexactly.pdf.

Bjorkholm:1998:ELS

[8] John E. Bjorkholm. EUV lithography — the successor to optical lithography? Intel Technology Journal, (Q3):8, 1998. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q31998/
articles/art_4.htm;   http://developer.intel.com/technology/
itj/q31998/pdf/euv.pdf.

Meieran:1998:CSM

[9] Eugene S. Meieran. 21st Century semiconductor manufacturing capabilities. Intel Technology Journal, (Q4):8, 1998. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q41998/
articles/art_1.htm;   http://developer.intel.com/technology/
itj/q41998/pdf/21manufacturing.pdf.

Brand:1998:IML

[10] A. Brand, A. Haranahalli, N. Hsieh, Y. C. Lin, G. Sery, N. Stenton, B. J. Woo, S. Ahmed, M. Bohr, S. Thompson, and S. Yang. Intel’s 0.25 micron, 2.0volts logic process technology. Intel Technology Journal, (Q3):9, 1998.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/q31998/articles/art_1.htm;   http://developer.intel.com/
technology/itj/q31998/pdf/p856.pdf.

ElGebaly:1998:CMS

[11] Hani ElGebaly. Characterization of multimedia streams of an H.323 terminal. Intel Technology Journal, (Q2):9, 1998. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q21998/
articles/art_5.htm;   http://developer.intel.com/technology/
itj/q21998/pdf/streams.pdf.

Hilton:1998:MOS

[12] Court Hilton. Manufacturing operations system design and analysis. Intel Technology Journal, (Q4):9, 1998. ISSN 1535-766X. URL http://

developer.intel.com/technology/itj/q41998/articles/art_3.htm;
 http://developer.intel.com/technology/itj/q41998/pdf/manf_design.
pdf.

Mitchell:1998:SPS

[13] Bradley Mitchell. Scalable platform services on the Intel TFLOPS supercomputer. Intel Technology Journal, (Q1):9, 1998. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q11998/
articles/art_4.htm;   http://developer.intel.com/technology/
itj/q11998/pdf/sps.pdf.

Cornea-Hasegan:1998:PIC

[14] Marius Cornea-Hasegan. Proving the IEEE correctness of iterative floating-point square root, divide, and remainder algorithms. Intel Technology Journal, (Q2):11, 1998. ISSN 1535-766X. URL http://

developer.intel.com/technology/itj/q21998/articles/art_3.htm;
 http://developer.intel.com/technology/itj/q21998/pdf/ieee.pdf.

Henry:1998:PIT

[15] Greg Henry, Pat Fay, Ben Cole, and Timothy G. Mattson. The performance of the Intel TFLOPS supercomputer. Intel Technology Journal, (Q1):11, 1998. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q11998/articles/art_2.htm;   http:/
/developer.intel.com/technology/itj/q11998/pdf/perf.pdf.

Hong:1998:OAF

[16] Yeoh Eng Hong, Lim Seong Leong, Wong Yik Choong, Lock Choon Hou, and Mahmud Adnan. An overview of advanced failure analysis techniques for Pentium and Pentium Pro microprocessors. Intel Technology Journal, (Q2):11, 1998. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q21998/articles/art_1.htm;   http:/
/developer.intel.com/technology/itj/q21998/pdf/fa.pdf.

Seshan:1998:QRI

[17] K. Seshan, T. J. Maloney, and K. J. Wu. Quality and reliability of Intel’s quarter micron process. Intel Technology Journal, (Q3):11, 1998.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/q31998/articles/art_2.htm;   http://developer.intel.com/
technology/itj/q31998/pdf/qa.pdf.

Toga:1998:DMC

[18] James Toga and Hani ElGebaly. Demystifying multimedia conferencing over the Internet using the H.323 set of standards. Intel Technology Journal, (Q2):11, 1998. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q21998/articles/art_4.htm;   http:/
/developer.intel.com/technology/itj/q21998/pdf/h323.pdf.

Garg:1998:ALS

[19] Sharad Garg, Robert Godley, Richard Griffiths, Andrew Pfiffer, Terry Prickett, David Robboy, Stan Smith, T. Mack Stallcup, and Stephan Zeisset. Achieving large scale parallelism through operating system resource management on the Intel TFLOPS supercomputer. Intel Technology Journal, (Q1):12, 1998. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q11998/articles/art_3.htm;   http:/
/developer.intel.com/technology/itj/q11998/pdf/tos.pdf.

Mattson:1998:OIT

[20] Timothy G. Mattson and Greg Henry. An overview of the Intel TFLOPS supercomputer. Intel Technology Journal, (Q1):12, 1998. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q11998/
articles/art_1.htm;   http://developer.intel.com/technology/
itj/q11998/pdf/overview.pdf.

Seligson:1998:PT

[21] Daniel Seligson. Planning for the 300mm transition. Intel Technology Journal, (Q4):14, 1998. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q41998/articles/art_4.htm;   http:/
/developer.intel.com/technology/itj/q41998/pdf/300mm.pdf.

Bradski:1998:CVF

[22] Gary R. Bradski. Computer vision face tracking for use in a perceptual user interface. Intel Technology Journal, (Q2):15, 1998. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q21998/
articles/art_2.htm;   http://developer.intel.com/technology/
itj/q21998/pdf/camshift.pdf.

Girkar:1998:IIM

[23] Milind Girkar, Mohammad R. Haghighat, Paul Grey, Hideki Saito, Nicholas Stavrakos, and Constantine D. Polychronopoulos. Illinois-Intel multithreading library: Multithreading support for Intel architecture based multiprocessor systems. Intel Technology Journal, (Q1):15, 1998.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/q11998/articles/art_5.htm;   http://developer.intel.com/
technology/itj/q11998/pdf/iml.pdf.

Thompson:1998:MST

[24] Scott Thompson, Paul Packan, and Mark Bohr. MOS scaling: Transistor challenges for the 21st Century. Intel Technology Journal, (Q3):19, 1998.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/q31998/articles/art_3.htm;   http://developer.intel.com/
technology/itj/q31998/pdf/trans.pdf.

Fischer:1999:PIP

[25] Stephen Fischer, James Mi, and Albert Teng. Pentium III processor serial number feature and applications. Intel Technology Journal, (Q2):6, May 17, 1999. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q21999/articles/art_3.htm;   http:/
/developer.intel.com/technology/itj/q21999/pdf/serial_number.
pdf.

Carver:1999:POS

[26] Kathy Carver, Chuck Fleckenstein, Joshua LeVasseur, and Stephan Zeisset. Porting operating system kernels to the IA-64 architecture for pre-silicon validation purposes. Intel Technology Journal, (Q4): 7, November 22, 1999. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q41999/articles/art_4.htm;   http:/
/developer.intel.com/technology/itj/q41999/pdf/porting.pdf.

Fleckenstein:1999:POS

[27] Chuck Fleckenstein, Kathy Carver, Joshua LeVasseur, and Stephan Zeisset. Porting operating system kernels to the IA-64 architecture for pre-silicon validation purposes. Intel Technology Journal, (Q4): 7, November 22, 1999. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q41999/articles/art_4.htm;   http:/
/developer.intel.com/technology/itj/q41999/pdf/porting.pdf.

Harrison:1999:CTF

[28] John Harrison, Ted Kubaska, Shane Story, and Ping Tak Peter Tang. The computation of transcendental functions on the IA-64 architecture. Intel Technology Journal, (Q4):7, November 22, 1999. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q41999/
articles/art_5.htm;   http://developer.intel.com/technology/
itj/q41999/pdf/transendental.pdf.

Yadav:1999:PFP

[29] Satyendra Yadav, Sanjay Bakshi, Dave Putzolu, and Raj Yavatkar. The Phoenix framework: A practical architecture for programmable networks. Intel Technology Journal, (Q3):7, August 4, 1999. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q31999/
articles/art_1.htm;   http://developer.intel.com/technology/
itj/q31999/pdf/phoenix.pdf.

Krishna:1999:CDE

[30] Bharat Krishna and Gil Kleinfeld. Circuit design environment and layout planning. Intel Technology Journal, (Q1):8, February 17, 1999. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q11999/
articles/art_2.htm;   http://developer.intel.com/technology/
itj/q11999/pdf/circuit_design.pdf.

Thakkar:1999:ISS

[31] Shreekant (Ticky) Thakkar and Tom Huff. The Internet Streaming SIMD Extensions. Intel Technology Journal, (Q2):8, May 17, 1999. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q21999/
articles/art_1.htm;   http://developer.intel.com/technology/
itj/q21999/pdf/simd_ext.pdf.

Dreke:1999:IPC

[32] Christian Dreke. Introduction to personal communication on the Internet. Intel Technology Journal, (Q3):9, August 4, 1999. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q31999/
articles/art_2.htm;   http://developer.intel.com/technology/
itj/q31999/pdf/per_comm.pdf.

Kumar:1999:SSH

[33] Vineet Kumar. Supplementary services in the H.323 IP multimedia telephony network. Intel Technology Journal, (Q3):9, August 4, 1999.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/q31999/articles/art_3.htm;   http://developer.intel.com/
technology/itj/q31999/pdf/telephony.pdf.

Nagbhushan:1999:NSA

[34] V. Nagbhushan, Yehuda Shiran, Satish Venkatesan, and Tamar Yehoshua. Nike’s software architecture and infrastructure: Enabling integrated solutions for gigahertz designs. Intel Technology Journal, (Q1):9, February 17, 1999. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q11999/articles/art_1.htm;   http:/
/developer.intel.com/technology/itj/q11999/pdf/nike.pdf.

ElGebaly:1999:RMR

[35] Hani ElGebaly. Reactive mechanisms for recovering audio performance in multimedia conferencing over packet switched networks. Intel Technology Journal, (Q3):10, August 4, 1999. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q31999/
articles/art_4.htm;   http://developer.intel.com/technology/
itj/q31999/pdf/audio.pdf.

Krishnapura:1999:CDF

[36] Shesha Krishnapura, Ty Tang, and Vipul Lal. CAD design flows development in a cross-platform computing environment. Intel Technology Journal, (Q1):10, February 17, 1999. ISSN 1535-766X. URL http://

developer.intel.com/technology/itj/q11999/articles/art_4.htm;
 http://developer.intel.com/technology/itj/q11999/pdf/cross_
platform.pdf.

OLeary:1999:FVI

[37] John O’Leary, Xudong Zhao, Rob Gerth, and Carl-Johan H. Seger. Formally verifying IEEE compliance of floating-point hardware. Intel Technology Journal, (Q1):10, February 17, 1999. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q11999/
articles/art_5.htm;   http://developer.intel.com/technology/
itj/q11999/pdf/floating_point.pdf.

Tal:1999:ALP

[38] Ady Tal, Vadim Bassin, Shay Gal-On, and Elena Demikhovsky. Assembly language programming tools for the IA-64 architecture. Intel Technology Journal, (Q4):10, November 22, 1999. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q41999/
articles/art_3.htm;   http://developer.intel.com/technology/
itj/q41999/pdf/assemble.pdf.

Keshava:1999:PIP

[39] Jagannath Keshava and Vladimir Pentkovski. Pentium III processor implementation tradeoffs. Intel Technology Journal, (Q2):11, May 17, 1999.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/q21999/articles/art_2.htm;   http://developer.intel.com/
technology/itj/q21999/pdf/impliment.pdf.

Wolf:1999:PMP

[40] Joe H. Wolf. Programming methods for the Pentium III processor’s streaming SIMD extensions using the VTune performance enhancement environment. Intel Technology Journal, (Q2):11, May 17, 1999. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q21999/
articles/art_6.htm;   http://developer.intel.com/technology/
itj/q21999/pdf/vtune.pdf.

Zagacki:1999:ASS

[41] Paul M. Zagacki, Deep Buch, Emile Hsieh, Daniel Melaku, Vladimir Pentkovski, and Hsien-Hsin Lee. Architecture of a 3D software stack for peak Pentium III processor performance. Intel Technology Journal, (Q2):11, May 17, 1999. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q21999/articles/art_4.htm;   http:/
/developer.intel.com/technology/itj/q21999/pdf/3d_stack.pdf.

Abel:1999:ATS

[42] James Abel, Kumar Balasubramanian, Mike Bargeron, Tom Craver, and Mike Phlipot. Applications tuning for streaming SIMD extensions. Intel Technology Journal, (Q2):13, May 17, 1999. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q21999/
articles/art_5.htm;   http://developer.intel.com/technology/
itj/q21999/pdf/apps_simd.pdf.

Sengupta:1999:DBT

[43] Sanjay Sengupta, Sandip Kundu, Sr.eejit Chakravarty, Praveen Parvathala, Rajesh Galivanche, George Kosonocky, Mike Rodgers, and TM Mak. Defect-based test: A key enabler for successful migration to structural test. Intel Technology Journal, (Q1):14, February 17, 1999.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/q11999/articles/art_6.htm;   http://developer.intel.com/
technology/itj/q11999/pdf/defect_based.pdf.

Uhlig:1999:SPS

[44] Richard Uhlig, Roman Fishtein, Oren Gershon, Israel Hirsh, and Hong Wang. SoftSDV: A pre-silicon software development environment for the IA-64 architecture. Intel Technology Journal, (Q4):14, November 22, 1999.

ISSN 1535-766X. URL http://developer.intel.com/technology/

itj/q41999/articles/art_2.htm;   http://developer.intel.com/
technology/itj/q41999/pdf/softsdv.pdf.

Dulong:1999:OII

[45] Carole Dulong, Rakesh Krishnaiyer, Dattatraya Kulkarni, Daniel Lavery, Wei Li, John Ng, and David Sehr. An overview of the Intel IA-64 compiler. Intel Technology Journal, (Q4):15, November 22, 1999. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q41999/
articles/art_1.htm;   http://developer.intel.com/technology/
itj/q41999/pdf/compiler.pdf.

Cornea-Hasegan:1999:IFP

[46] Marius Cornea-Hasegan and Bob Norin. IA-64 floating-point operations and the IEEE standard for binary floating-point arithmetic. Intel Technology Journal, (Q4):16, November 22, 1999. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q41999/
articles/art_6.htm;   http://developer.intel.com/technology/
itj/q41999/pdf/ia64fpbf.pdf;   http://gec.di.uminho.pt/discip/
minf/ac0203/icca03/ia64fpbf1.pdf.

Chan:1999:CCD

[47] Tim Chan, Amit Chowdhary, Bharat Krishna, Artour Levin, Gary Meeker, and Naresh Sehgal. Challenges of CAD development for datapath design. Intel Technology Journal, (Q1):17, February 17, 1999. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q11999/
articles/art_3.htm;   http://developer.intel.com/technology/
itj/q11999/pdf/datapath_design.pdf.

Anonymous:1999:SII

[48] Anonymous. Special issue on Intel Architecture 64-bit (IA-64) technology. Intel Technology Journal, (Q4), November 22, 1999. ISSN 1535-766X.

URL http://developer.intel.com/technology/itj/q41999.htm.

Weatherspoon:2000:OIS

[49] Sultan Weatherspoon. Overview of IEEE 802.11b security. Intel Technology Journal, (Q2):5, May 5, 2000. ISSN 1535-766X. URL http://

developer.intel.com/technology/itj/q22000/articles/art_5.htm;
 http://developer.intel.com/technology/itj/q22000/pdf/art_5.
pdf.

Ames:2000:ETG

[50] Phillip Ames and John Gabor. The evolution of third-generation cellular standards. Intel Technology Journal, (Q2):6, May 5, 2000. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q22000/
articles/art_6.htm;   http://developer.intel.com/technology/
itj/q22000/pdf/art_6.pdf.

Coyne:2000:BAM

[51] Brad Coyne and Sonja K. Sandeen. e-Business asset management and capacity planning. Intel Technology Journal, (Q4):6, 2000. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q42000/
articles/art_4.htm;   http://developer.intel.com/technology/
itj/q42000/pdf/assetmngt.pdf.

Schlierf:2000:CSR

[52] Michael Schlierf, Michael Fuller, and Mike Kerr. Certifying service reliability in data centers. Intel Technology Journal, (Q4):6, 2000. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q42000/
articles/art_3.htm;   http://developer.intel.com/technology/
itj/q42000/pdf/certify.pdf.

Aneja:2000:CPF

[53] Atul Aneja, Chia Rowan, and Brian Brooksby. Corporate portal framework for transforming content chaos on intranets. Intel Technology Journal, (Q1):7, February 3, 2000. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q12000/articles/art_3.htm;   http:/
/developer.intel.com/technology/itj/q12000/pdf/portal.pdf.

Dory:2000:SCJ

[54] Tom Dory, Kenji Takahashi, Tomomi Kume, Jiro Kubota, Seiichiro Seki, and Takaharu Fujiyama. Simultaneous chip-join and underfill assembly technology for flip-chip packaging. Intel Technology Journal, (Q3):7, August 22, 2000. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q32000/articles/art_5.htm;   http:/
/developer.intel.com/technology/itj/q32000/pdf/assembly.pdf.

Fleming:2000:EAA

[55] Kristoffer Fleming, Robert J. Hunter, Jon Inouye, and Jeffrey Schiffer. Enabling always on, always connected (AOAC) computing with Bluetooth technology. Intel Technology Journal, (Q2):7, May 5, 2000. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q22000/
articles/art_3.htm;   http://developer.intel.com/technology/
itj/q22000/pdf/art_3.pdf.

Kardach:2000:BAO

[56] James Kardach. Bluetooth architecture overview. Intel Technology Journal, (Q2):7, May 5, 2000. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q22000/articles/art_1.htm;   http:/
/developer.intel.com/technology/itj/q22000/pdf/art_1.pdf.

Kirby:2000:IBT

[57] Graham Kirby. Integrating Bluetooth technology into mobile products. Intel Technology Journal, (Q2):8, May 5, 2000. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q22000/
articles/art_4.htm;   http://developer.intel.com/technology/
itj/q22000/pdf/art_4.pdf.

Mencinger:2000:MBM

[58] Nicholas P. Mencinger. A mechanism-based methodology for processor package reliability assessments. Intel Technology Journal, (Q3): 8, August 22, 2000. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q32000/articles/art_6.htm;   http:/
/developer.intel.com/technology/itj/q32000/pdf/reliability.pdf.

Tadayon:2000:TCD

[59] Pooya Tadayon. Thermal challenges during microprocessor testing. Intel Technology Journal, (Q3):8, August 22, 2000. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q32000/
articles/art_7.htm;   http://developer.intel.com/technology/
itj/q32000/pdf/thermal.pdf.

Hodgson:2000:IEE

[60] Alan Hodgson. Intel eBusiness engineering release management and application landing. Intel Technology Journal, (Q4):9, 2000. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q42000/
articles/art_5.htm;   http://developer.intel.com/technology/
itj/q42000/pdf/releasemngt.pdf.

Lii:2000:FCT

[61] Mirng-Ji Lii, Bob Sankman, Hamid Azimi, Hwai Peng Yeoh, and Yuejin Guo. Flip-chip technology on organic pin grid array packages. Intel Technology Journal, (Q3):9, August 22, 2000. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q32000/
articles/art_2.htm;   http://developer.intel.com/technology/
itj/q32000/pdf/flipchip.pdf.

Fleming:2000:AOI

[62] Kris Fleming, Uma Gadamsetty, Robert J. Hunter, Sr.ikanth Kambhatla, Sr.idhar Rajagopal, and Sundaram Ramakesavan. Architectural overview of Intel’s Bluetooth software stack. Intel Technology Journal, (Q2):10, May 5, 2000. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q22000/articles/art_2.htm;   http:/
/developer.intel.com/technology/itj/q22000/pdf/art_2.pdf.

Mahajan:2000:EMP

[63] Ravi Mahajan, Ken Brown, and Vasu Atluri. The evolution of microprocessor packaging. Intel Technology Journal, (Q3):10, August 22, 2000.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/q32000/articles/art_1.htm;   http://developer.intel.com/
technology/itj/q32000/pdf/package.pdf.

Bickerstaff:2000:IIC

[64] Cindy Bickerstaff, Sally Hambridge, Lynne Marchi, Tod Oace, Stacy Purcell, Jeff Sedayao, Charles Smothers, and Ken True. Intel’s Internet connectivity: Evolution, technical architecture, and future directions. Intel Technology Journal, (Q1):11, February 3, 2000. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q12000/
articles/art_1.htm;   http://developer.intel.com/technology/
itj/q12000/pdf/internet.pdf.

Fieldhouse:2000:EIB

[65] Tom Fieldhouse, Mark Greinke, Jay Hahn-Steichen, Jackson He, and John Vliet. Evolution of Intel’s e-business data center architecture from yesterday to tomorrow. Intel Technology Journal, (Q4):11, 2000. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q42000/
articles/art_1.htm;   http://developer.intel.com/technology/
itj/q42000/pdf/datacenter.pdf.

Marchi:2000:IAS

[66] Lynne Marchi, Sridhar Mahankali, and Jeff Sedayao. IP addressing space design issues for Internet data centers. Intel Technology Journal, (Q4):12, 2000. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q42000/articles/art_2.htm;   http:/
/developer.intel.com/technology/itj/q42000/pdf/ipaddress.pdf.

Calame:2000:CDS

[67] Paul Calame, Ravi Nannapaneni, Scott Peterson, Jay Turpin, and James Yu. Cockpit: Decision support tool for factory operations and supply chain management. Intel Technology Journal, (Q1):13, February 3, 2000.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/q12000/articles/art_2.htm;   http://developer.intel.com/
technology/itj/q12000/pdf/cockpit.pdf.

Vicente:2000:MEN

[68] John Vicente, Harold Cartmill, Glen Maxson, Shelby Siegel, and Russ Fenger. Managing enhanced network services: A pragmatic view of policybased management. Intel Technology Journal, (Q1):13, February 3, 2000.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/q12000/articles/art_5.htm;   http://developer.intel.com/
technology/itj/q12000/pdf/policybased.pdf.

OSullivan:2000:IIB

[69] Patricia J. O’Sullivan and Don S. Whitecar. Implementing an industry e-business initiative: Getting to RosettaNet. Intel Technology Journal, (Q1):15, February 3, 2000. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q12000/articles/art_4.htm;   http:/
/developer.intel.com/technology/itj/q12000/pdf/rosettanet.pdf.

Viswanath:2000:TPC

[70] Ram Viswanath, Vijay Wakharkar, Abhay Watwe, and Vassou Lebonheur. Thermal performance challenges from silicon to systems. Intel Technology Journal, (Q3):16, August 22, 2000. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q32000/
articles/art_4.htm;   http://developer.intel.com/technology/
itj/q32000/pdf/thermal_perf.pdf.

Polka:2000:PLI

[71] Lesley Polka, Shamala Chickamenahalli, Chee-Yee Chung, David G. Figueroa, Yuan-Liang Li, Kim Merley, Dustin Wood, and Larry Zu. Package-level interconnect design for optimum electrical performance. Intel Technology Journal, (Q3):17, August 22, 2000. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q32000/
articles/art_3.htm;   http://developer.intel.com/technology/
itj/q32000/pdf/interconnect.pdf.

Bentley:2001:VIP

[72] Bob Bentley and Rand Gray. Validating the Intel Pentium 4 processor. Intel Technology Journal, (Q1):8, February 12, 2001. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q12001/
articles/art_3.htm;   http://developer.intel.com/technology/
itj/q12001/pdf/art_3.pdf.

Bik:2001:EEP

[73] Aart Bik, Milind Girkar, Paul Grey, and Xinmin Tian. Efficient exploitation of parallelism on Pentium III and Pentium 4 processor-based systems. Intel Technology Journal, (Q1):9, February 12, 2001. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q12001/
articles/art_6.htm;   http://developer.intel.com/technology/
itj/q12001/pdf/art_6.pdf.

Gunther:2001:MII

[74] Stephen H. Gunther, Frank Binns, Douglas M. Carmean, and Jonathan C. Hall. Managing the impact of increasing microprocessor power consumption. Intel Technology Journal, (Q1):9, February 12, 2001. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q12001/
articles/art_4.htm;   http://developer.intel.com/technology/
itj/q12001/pdf/art_4.pdf.

Hinton:2001:MPP

[75] Glenn Hinton, Dave Sager, Mike Upton, Darrell Boggs, Doug Carmean, Alan Kyker, and Patrice Roussel. The microarchitecture of the Pentium 4 processor. Intel Technology Journal, (Q1):13, February 12, 2001.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/q12001/articles/art_2.htm;   http://developer.intel.com/
technology/itj/q12001/pdf/art_2.pdf.

Kumar:2001:INI

[76] Rajesh Kumar. Interconnect and noise immunity design for the Pentium 4 processor. Intel Technology Journal, (Q1):12, February 12, 2001.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/q12001/articles/art_5.htm;   http://developer.intel.com/
technology/itj/q12001/pdf/art_5.pdf.

Tetrick:2001:DPP

[77] R. Scott Tetrick, Blaise Fanning, Robert Greiner, Tom Huff, Lance Hacking, David Hill, Srinivas Chennupaty, David Koufaty, Subba Palacharla, Jeff Rabe, and Mike Derr. A discussion of PC platform balance: the Intel Pentium 4 processor-based platform solutions. Intel Technology Journal, (Q1):13, February 12, 2001. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q12001/articles/art_7.htm;   http:/
/developer.intel.com/technology/itj/q12001/pdf/art_7.pdf.

Volk:2001:REC

[78] Andrew M. Volk, Peter A. Stoll, and Paul Metrovich. Recollections of early chip development at Intel. Intel Technology Journal, (Q1):

12, February 12, 2001. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q12001/articles/art_1.htm;   http:/
/developer.intel.com/technology/itj/q12001/pdf/art_1.pdf.

Bard:2001:WCP

[79] Steve Bard. Wireless convergence of PC and consumer electronics in the e-home. Intel Technology Journal, (Q2):11, May 10, 2001. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q22001/
articles/art_2.htm;   http://developer.intel.com/technology/
itj/q22001/pdf/art_2.pdf.

Chou:2001:OCA

[80] Joey Chou. Open CPE architecture: A solution to the delivery of integrated services over broadband. Intel Technology Journal, (Q2):10, May 10, 2001. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q22001/articles/art_1.htm;   http:/
/developer.intel.com/technology/itj/q22001/pdf/art_1.pdf.

Foerster:2001:UWT

[81] Jeff Foerster, Evan Green, Srinivasa Somayazulu, and David Leeper. Ultra-wideband technology for short- or medium-range wireless communications. Intel Technology Journal, (Q2):11, May 10, 2001. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q22001/
articles/art_4.htm;   http://developer.intel.com/technology/
itj/q22001/pdf/art_4.pdf.

Merrick:2001:EPH

[82] Craig Merrick, Robert Dunstan, and Michael Jeronimo. Extending the PC in the home. Intel Technology Journal, (Q2):13, May 10, 2001. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q22001/
articles/art_3.htm;   http://developer.intel.com/technology/
itj/q22001/pdf/art_3.pdf.

Regnier:2001:CSL

[83] Greg Regnier. CSP: A system-level architecture for scalable communication services. Intel Technology Journal, (Q2):6, May 10, 2001. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q22001/
articles/art_5.htm;   http://developer.intel.com/technology/
itj/q22001/pdf/art_5.pdf.

Bell:2001:LAA

[84] Genevieve Bell. Looking across the Atlantic: Using ethnographic methods to make sense of Europe. Intel Technology Journal, (Q3):

10, August 1, 2001. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q32001/articles/art_1.htm;   http:/
/developer.intel.com/technology/itj/q32001/pdf/art_1.pdf.

Dulong:2001:MCI

[85] Carole Dulong, Priti Shrivastav, and Azita Refah. The making of a compiler for the Intel(R) Itanium processor. Intel Technology Journal, (Q3):7, August 1, 2001. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q32001/articles/art_4.htm;   http:/
/developer.intel.com/technology/itj/q32001/pdf/art_4.pdf.

Rusch:2001:IWC

[86] Leslie Ann Rusch. Indoor wireless communications: Capacity and coexistence on the unlicensed bands. Intel Technology Journal, (Q3): 10, August 1, 2001. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q32001/articles/art_2.htm;   http:/
/developer.intel.com/technology/itj/q32001/pdf/art_2.pdf.

Srinivasan:2001:ERA

[87] Kumud M. Srinivasan. The evolving role of automation in Intel(R) microprocessor development and manufacturing. Intel Technology Journal, (Q3):7, August 1, 2001. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q32001/articles/art_3.htm;   http:/
/developer.intel.com/technology/itj/q32001/pdf/art_3.pdf.

Boss:2001:ASI

[88] Scott Boss. Audio software for the Intel(R) Play Computer Sound Morpher. Intel Technology Journal, (Q4):11, November 17, 2001. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q42001/
articles/art_5.htm;   http://developer.intel.com/technology/
itj/q42001/pdf/Audio Software for the   Intel Play Computer Sound
Morpher.pdf.

Boss:2001:DST

[89] Scott Boss, Henry Bruce, Charlie Case, and Kendal Miller. Developing smart toys—from idea to product. Intel Technology Journal, (Q4): 9, November 17, 2001. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q42001/articles/art_2.htm;   http:/
/developer.intel.com/technology/itj/q42001/pdf/Developing Smart
 Toys-From Idea to Product.pdf.

Conti:2001:TPP

[90] Michael M. Conti, Riley Jackson, and Henry Bruce. Technology and play pattern: Intel(R) Play Digital Movie Creator. Intel Technology Journal, (Q4):7, November 17, 2001. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q42001/
articles/art_6.htm;   http://developer.intel.com/technology/
itj/q42001/pdf/Technology and Play   Pattern.pdf.

DHooge:2001:GDP

[91] Herman D’Hooge and Michael Goldsmith. Game design principles for the Intel(R) Play Me2Cam* virtual game system. Intel Technology Journal, (Q4):9, November 17, 2001. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/q42001/
articles/art_4.htm;   http://developer.intel.com/technology/
itj/q42001/pdf/Game Design Principles   for the Intel Play Me2Cam
Virtual Game System.pdf.

DHooge:2001:HST

[92] Herman D’Hooge and Melanie Goldstein. History of the Smart Toy Lab and Intel(R) Play Toys. Intel Technology Journal, (Q4): 6, November 17, 2001. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q42001/articles/art_1.htm;   http:/
/developer.intel.com/technology/itj/q42001/pdf/History of the
Smart Toy   Lab and Intel Play Toys.pdf.

Jelinek:2001:DIP

[93] Lenka Jelinek, Geoff Peters, Jim Okuley, and Steve McGowan. Dissection of the Intel(R) Play QX3 computer microscope. Intel Technology Journal, (Q4):10, November 17, 2001. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/q42001/articles/art_3.htm;   http:/
/developer.intel.com/technology/itj/q42001/pdf/Dissection of
the Intel   Play QX3 Computer Microscope.pdf.

Chao:2002:PQ

[94] Lin Chao. Preface: Q1. 2002. Intel Technology Journal, 6(1):2, Febru-

ary 2002. ISSN 1535-766X. URL http://developer.intel.com/
technology/itj/2002/volume06issue01/vol6iss1_hyper_threading_
technology.pdf.

Cross:2002:IFQ

[95] Robert L. Cross. ITJ foreword Q1, 2002: Intel(R) hyper-threading technology. Intel Technology Journal, 6(1):3, February 2002. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2002/
volume06issue01/vol6iss1_hyper_threading_technology.pdf.

Marr:2002:HTT

[96] Deborah T. Marr, Frank Binns, David L. Hill, Glenn Hinton, David A. Koufaty, J. Alan Miller, and Michael Upton. Hyper-threading technology architecture and microarchitecture. Intel Technology Journal, 6(1):4–15, February 2002. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2002/volume06issue01/vol6iss1_hyper_threading_
technology.pdf.

Burns:2002:PSV

[97] David Burns. Pre-silicon validation of hyper-threading technology. Intel Technology Journal, 6(1):16–21, February 2002. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2002/
volume06issue01/vol6iss1_hyper_threading_technology.pdf.

Wang:2002:SPE

[98] Hong Wang, Perry H. Wang, Ross Dave Weldon, Scott M. Ettinger, Hideki Saito, Milind Girkar, Steve Shih wei Liao, and John P. Shen. Speculative precomputation: Exploring the use of multithreading for latency tools. Intel Technology Journal, 6(1):22–35, February 2002. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2002/
volume06issue01/vol6iss1_hyper_threading_technology.pdf.

Tian:2002:IOC

[99] Xinmin Tian, Aart Bik, Milind Girkar, Paul Grey, Hideki Saito, and Ernesto Su. Intel(R) OpenMP C++/Fortran compiler for hyper-threading technology: Implementation and performance. Intel Technology Journal, 6(1):36–46, February 2002. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2002/
volume06issue01/vol6iss1_hyper_threading_technology.pdf.

Chen:2002:MAH

[100] Yen-Kuang Chen, Matthew Holliman, Eric Debes, Sergey Zheltov, Alexander Knyazev, Stanislav Bratanov, Roman Belenov, and Ishmael Santos. Media applications on hyper-threading technology. Intel Technology Journal, 6(1):47–57, February 2002. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2002/
volume06issue01/vol6iss1_hyper_threading_technology.pdf.

Magro:2002:HTT

[101] William Magro, Paul Petersen, and Sanjiv Shah. Hyper-threading technology: Impact on compute-intensive workloads. Intel Technology Journal, 6(1):58–66, February 2002. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2002/volume06issue01/vol6iss1_hyper_
threading_technology.pdf.

Anonymous:2002:P

[102] Anonymous. Preface. Intel Technology Journal, 6(2):3, May 16, 2002.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2002/volume06issue02/vol6iss2_semiconductor_technology_and_
manufacturing.pdf.

Bohr:2002:STM

[103] Mark Bohr. Semiconductor technology and manufacturing—current and future. Intel Technology Journal, 6(2):4, May 16, 2002. ISSN

1535-766X.  URL http://developer.intel.com/technology/itj/
2002/volume06issue02/foreword.htm;   http://developer.intel.
com/technology/itj/2002/volume06issue02/bohr_bio.htm.

Thompson:2002:LTF

[104] Scott Thompson, Mohsen Alavi, Makarem Hussein, Pauline Jacob, Chris Kenyon, Peter Moon, Matthew Prince, Sam Sivakumar, Sunit Tyagi, Jr., and Mark Bohr. 130nm logic technology featuring 60nm transistors, low-K dielectrics and Cu interconnects. Intel Technology Journal, 6(2):5–13, May 16, 2002. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2002/
volume06issue02/art01_130nmlogic/p01_abstract.htm;   http://
developer.intel.com/technology/itj/2002/volume06issue02/art01_
130nmlogic/vol6iss2_art01.pdf.

Natarajan:2002:PDM

[105] Sanjay Natarajan, Melton Bost, Derek Fisher, David Krick, Chris Kenyon, Chris Kardas, Chris Parker, and Robert Gasser, Jr. Process development and manufacturing of high-performance microprocessors on 300mm wafers. Intel Technology Journal, 6(2):14–22, May 16, 2002. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2002/
volume06issue02/art02_processdev/p01_abstract.htm;   http://
developer.intel.com/technology/itj/2002/volume06issue02/art02_
processdev/vol6iss2_art02.pdf.

Fazio:2002:EFM

[106] Al Fazio, Stephen Keeney, and Stefan Lai. ETOX^{T M} flash memory technology: Scaling and integration challenges. Intel Technology Journal, 6(2): 23–30, May 16, 2002. ISSN 1535-766X. URL http://developer.intel.

com/technology/itj/2002/volume06issue02/art03_flashmemory/p01_
abstract.htm;   http://developer.intel.com/technology/itj/2002/
volume06issue02/art03_flashmemory/vol6iss2_art03.pdf.

Kuhn:2002:IMS

[107] Kelin J. Kuhn, Shahriar Ahmed, Peter Vandervoorn, Anand Murthy, Borna Obradovic, Kartik Raol, Wei kai Shih, Iwen Chao, Ian Post, and Steve Chambers. Integration of mixed-signal elements into a highperformance digital CMOS process. Intel Technology Journal, 6(2):31–41, May 16, 2002. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2002/volume06issue02/art04_mixedsignalelements/
p01_abstract.htm;   http://developer.intel.com/technology/itj/
2002/volume06issue02/art04_mixedsignalelements/vol6iss2_art04.
pdf.

Doyle:2002:TEP

[108] Brian Doyle, Reza Arghavani, Doug Barlage, Suman Datta, Mark Doczy, Jack Kavalieros, Anand Murthy, and Robert Chau. Transistor elements for 30nm physical gate length and beyond. Intel Technology Journal, 6(2):42–54, May 16, 2002. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2002/
volume06issue02/art05_transistorarch/p01_abstract.htm;   http:
//developer.intel.com/technology/itj/2002/volume06issue02/art05_
transistorarch/vol6iss2_art05.pdf.

Silverman:2002:ILR

[109] Peter J. Silverman. The Intel lithography roadmap. Intel Technology Journal, 6(2):55–61, May 16, 2002. ISSN 1535-766X. URL http://

developer.intel.com/technology/itj/2002/volume06issue02/art06_
lithographyroadmap/p01_abstract.htm;   http://developer.intel.
com/technology/itj/2002/volume06issue02/art06_lithographyroadmap/
vol6iss2_art06.pdf.

Mahajan:2002:EDP

[110] Ravi Mahajan, Raj Nair, Vijay Wakharkar, Johanna Swan, John Tang, and Gilroy Vandentop. Emerging directions for packaging technologies. Intel Technology Journal, 6(2):62–75, May 16, 2002. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2002/
volume06issue02/art07_emergingdirections/p01_abstract.htm;

http://developer.intel.com/technology/itj/2002/volume06issue02/
art07_emergingdirections/vol6iss2_art07.pdf.

Chao:2002:P

[111] Lin Chao. Preface. Intel Technology Journal, 6(3):3, August 15, 2002.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2002/volume06issue03/vol6iss3_networkprocessors.pdf.

Finnegan:2002:FQI

[112] Jim Finnegan. Foreword Q3, 2002 ITJ: Innovations in network processors. Intel Technology Journal, 6(3):4–5, August 15, 2002.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2002/volume06issue03/vol6iss3_networkprocessors.pdf.

Adiletta:2002:NGI

[113] Matthew Adiletta, Mark Rosenbluth, Debra Bernstein, Gilbert Wolrich, and Hugh Wilkinson. The next generation of Intel IXP network processors. Intel Technology Journal, 6(3):6–18, August 15, 2002. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2002/
volume06issue03/art01_nextgenixp/p01_abstract.htm;   http://
developer.intel.com/technology/itj/2002/volume06issue03/art01_
nextgenixp/vol6iss3_art01.pdf.

Lakshmanamurthy:2002:NPP

[114] Sridhar Lakshmanamurthy, Kin-Yip Liu, Yim Pun, Larry Huston, and Uday Naik. Network processor performance analysis methodology. Intel Technology Journal, 6(3):19–28, August 15, 2002. ISSN

1535-766X.  URL http://developer.intel.com/technology/itj/
2002/volume06issue03/art04_performanceanalysis/p01_abstract.
htm;   http://developer.intel.com/technology/itj/2002/volume06issue03/
art04_performanceanalysis/vol6iss3_art04.pdf.

Adiletta:2002:PSA

[115] Matthew Adiletta, Donald Hooper, and Myles Wilde. Packet over SONET: Achieving 10 Gigabit/sec packet processing with an IXP2800. Intel Technology Journal, 6(3):29–39, August 15, 2002. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2002/
volume06issue03/art05_packetoversonet/p01_abstract.htm;   http:
//developer.intel.com/technology/itj/2002/volume06issue03/art05_
packetoversonet/vol6iss3_art05.pdf.

Feghali:2002:SAP

[116] Wajdi Feghali, Brad Burres, Gilbert Wolrich, and Douglas Carrigan. Security: Adding protection to the network via the network processor. Intel Technology Journal, 6(3):40–49, August 15, 2002. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2002/
volume06issue03/art02_security/p01_abstract.htm;   http://developer.
intel.com/technology/itj/2002/volume06issue03/art02_security/
vol6iss3_art02.pdf.

Naik:2002:IPF

[117] Uday Naik, Alex Shoykhet, Larry Huston, Donald Hooper, Raj Yavatkar, Duke Tallam, Travis Schluessler, Prashant Chandra, and Adrian Georgescu. IXA portability framework: Preserving software investment in network processor applications. Intel Technology Journal, 6(3):50–60, August 15, 2002. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2002/volume06issue03/art06_ixaportability/p01_
abstract.htm;   http://developer.intel.com/technology/itj/2002/
volume06issue03/art06_ixaportability/vol6iss3_art06.pdf.

Vipat:2002:NPB

[118] Harsh Vipat, Philip Mathew, Manohar Ruben Castelino, and Auro Tripathy. Network processor building blocks for all-IP wireless network. Intel Technology Journal, 6(3):61–69, August 15, 2002. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2002/
volume06issue03/art08_processorbuildingblocks/p01_abstract.htm;
 http://developer.intel.com/technology/itj/2002/volume06issue03/
art08_processorbuildingblocks/vol6iss3_art08.pdf.

Sydir:2002:IVA

[119] Jaroslaw Sydir, Prashant Chandra, Alok Kumar, Sridhar Lakshmanamurthy, Longsong Lin, and Muthaiah Venkatachalam. Implementing voice over AAL2 on a network processor. Intel Technology Journal, 6(3):70–82, August 15, 2002. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2002/volume06issue03/art07_voiceoveraal2/p01_
abstract.htm;   http://developer.intel.com/technology/itj/2002/
volume06issue03/art07_voiceoveraal2/vol6iss3_art07.pdf.

Bhamidipati:2002:CMI

[120] Ram Bhamidipati, Ahmad Zaidi, Siva Makineni, Kah K. Low, Robert Chen, Kin-Yip Liu, and Jack Dahlgren. Challenges and methodologies for implementing high-performance network processors. Intel Technology Journal, 6(3):83–92, August 15, 2002. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2002/
volume06issue03/art03_challenges/p01_abstract.htm;   http://
developer.intel.com/technology/itj/2002/volume06issue03/art03_
challenges/vol6iss3_art03.pdf.

Chao:2002:DCD

[121] Lin Chao. Digital content distribution in the home. Intel Technology Journal, 6(4):3, November 15, 2002. ISSN 1535-766X. URL http://

developer.intel.com/technology/itj/2002/volume06issue04/vol6iss4_
interoperable_home_infrastructure.pdf.

Weinrib:2002:IHI

[122] Abel Weinrib and Gerald Holzhammer. Interoperable home infrastructure. Intel Technology Journal, 6(4):4, November 15, 2002. ISSN

1535-766X.  URL http://developer.intel.com/technology/itj/
2002/volume06issue04/foreword.htm;   http://developer.intel.
com/technology/itj/2002/volume06issue04/weinrib_bio.htm;   http:
//developer.intel.com/technology/itj/2002/volume06issue04/holzhammer_
bio.htm.

Rasheed:2002:HIF

[123] Yasser Rasheed, Jim Edwards, and Charlie Tai. Home interoperability framework for the digital home. Intel Technology Journal, 6(4):5– 16, November 15, 2002. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2002/volume06issue04/art01_interop/
p01_abstract.htm;   http://developer.intel.com/technology/itj/
2002/volume06issue04/art01_interop/vol6iss4_art01.pdf.

Rasheed:2002:HQM

[124] Yasser Rasheed and John Ritchie. High-quality media distribution in the home. Intel Technology Journal, 6(4):17–29, November 15, 2002. ISSN

1535-766X.  URL http://developer.intel.com/technology/itj/
2002/volume06issue04/art02_distribution/p01_abstract.htm;   http:
//developer.intel.com/technology/itj/2002/volume06issue04/art02_
distribution/vol6iss4_art02.pdf.

Walker:2002:RFE

[125] Mark R. Walker, Jim Edwards, Michael Jeronimo, John G. Ritchie, and Ylian Saint-Hilaire. Remote I/O: Freeing the experience from the platform with UPnP* architecture. Intel Technology Journal, 6(4):30–36, November 15, 2002. ISSN 1535-766X. URL http://developer.intel.

com/technology/itj/2002/volume06issue04/art03_remote_io/p01_
abstract.htm;   http://developer.intel.com/technology/itj/2002/
volume06issue04/art03_remote_io/vol6iss4_art03.pdf.

Ellison:2002:HNS

[126] Carl M. Ellison. Home network security. Intel Technology Journal, 6(4): 37–48, November 15, 2002. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2002/volume06issue04/art04_security/
p01_abstract.htm;   http://developer.intel.com/technology/itj/
2002/volume06issue04/art04_security/vol6iss4_art04.pdf.

Ripley:2002:CPD

[127] Michael Ripley, C. Brendan S. Traw, Steve Balogh, and Michael Reed. Content protection in the digital home. Intel Technology Journal, 6(4): 49–55, November 15, 2002. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2002/volume06issue04/art05_protection/
p01_abstract.htm;   http://developer.intel.com/technology/itj/
2002/volume06issue04/art05_protection/vol6iss4_art05.pdf.

Krishnamurthy:2002:MDD

[128] Lakshman Krishnamurthy, Steven Conner, Mark Yarvis, Jasmeet Chhabra, Carl Ellison, Chuck Brabenac, and Ernest Tsui. Meeting the demands of the digital home with high-speed multi-hop wireless networks. Intel Technology Journal, 6(4):57–68, November 15, 2002. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2002/
volume06issue04/art06_wireless/p01_abstract.htm;   http://developer.
intel.com/technology/itj/2002/volume06issue04/art06_wireless/
vol6iss4_art06.pdf.

Gokulrangan:2002:IUI

[129] Venkat R. Gokulrangan. Internetworking using IPv6 technology inside and outside the home. Intel Technology Journal, 6(4):69–77, November 15, 2002. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2002/volume06issue04/art07_ipv6/p01_abstract.
htm;   http://developer.intel.com/technology/itj/2002/volume06issue04/
art07_ipv6/vol6iss4_art07.pdf2.

Chao:2003:PQIa

[130] Lin Chao. Preface Q1, 2003, ITJ. Intel Technology Journal, 7(1):3, February 2003. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2003/volume07issue01/foreword.htm.

Rattner:2003:FMR

[131] Justin Rattner. Foreword: Managed runtime technologies. Intel Technology Journal, 7(1):4, February 2003. ISSN 1535-766X. URL http://

developer.intel.com/technology/itj/2003/volume07issue01/foreword.
htm.

Cierniak:2003:ORP

[132] Michal Cierniak, Marsha Eng, Neal Glew, Brian Lewis, and James Stichnoth. The Open Runtime Platform: a flexible high-performance managed runtime environment. Intel Technology Journal, 7(1):5–18, February 2003.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2003/volume07issue01/art01_orp/p01_abstract.htm.

Adl-Tabatabai:2003:SDC

[133] Ali-Reza Adl-Tabatabai, Jay Bharadwaj, Dong-Yuan Chen, Anwar Ghuloum, Vijay Menon, Brian Murphy, Mauricio Serrano, and Tatiana Shpeisman. The StarJIT compiler: a dynamic compiler for managed runtime environments. Intel Technology Journal, 7(1):19–31, February 2003.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2003/volume07issue01/art02_starjit/p01_abstract.htm.

Chow:2003:EJP

[134] Kingsum Chow, Ricardo Morin, and Kumar Shiv. Enterprise Java performance: Best practices. Intel Technology Journal, 7(1):32–46, February 2003. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2003/volume07issue01/art03_java/p01_abstract.
htm.

Vorobiov:2003:DOW

[135] George Vorobiov, Carl Dichter, John Benninghoff, and Charlie Hewett. Developing and optimizing Web applications on ASP .NET platform. Intel Technology Journal, 7(1):47–59, February 2003. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2003/
volume07issue01/art04_optimizing/p01_abstract.htm.

Aissi:2003:RAW

[136] Selim Aissi. Runtime environment security models. Intel Technology Journal, 7(1):60–67, February 2003. ISSN 1535-766X. URL http://

developer.intel.com/technology/itj/2003/volume07issue01/art05_
security/p01_abstract.htm.

Comp:2003:RAW

[137] Lynn Comp and Tim Dobbing. Runtime abstractions in the wireless and handheld space. Intel Technology Journal, 7(1):68–76, February 2003.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2003/volume07issue01/art07_mobile/p01_abstract.htm.

Drews:2003:MRE

[138] Paul Drews, Doug Sommer, Roger Chandler, and Terry Smith. Managed runtime environments for next-generation mobile devices. Intel Technology Journal, 7(1):77–82, February 2003. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2003/
volume07issue01/art07_mobile/p01_abstract.htm.

Chao:2003:PQIb

[139] Lin Chao. Preface Q2, 2003 ITJ. Intel Technology Journal, 7(3):3, February 2003. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2003/volume07issue02/vol7iss2.pdf.

Chandrasekher:2003:FIC

[140] Anand Chandrasekher and David Perlmutter. Foreword: Intel(R) Centrino^{T M} mobile technology. Intel Technology Journal, 7(2):4–5,

May 2003. ISSN 1535-766X. URL http://developer.intel.com/
technology/itj/2003/volume07issue02/foreword.htm.

Chinn:2003:MPP

[141] Gordon Chinn, Sanjiv Desai, Eric DiStefano, Krishnan Ravichandran, and Shreekant (Ticky) Thakkar. Mobile PC platforms enabled with Intel(R) Centrino^{T M} mobile technology. Intel Technology Journal, 7(2): 6–15, May 2003. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2003/volume07issue02/art01_centrino/p01_abstract.
htm.

Trivedi:2003:IBL

[142] Shailesh Trivedi, Tom Shewchuk, Aditya Sreenivas, Terry Fletcher, and Michael de la Cruz. Innovation brings low power integrated graphics to the Intel(R) Centrino^{T M} mobile technology platform. Intel Technology Journal, 7(2):16–20, May 2003. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2003/volume07issue02/art02_graphics/
p01_abstract.htm.

Gochman:2003:IPM

[143] Simcha Gochman, Ronny Ronen, Ittai Anati, Ariel Berkovits, Tsvika Kurts, Alon Naveh, Ali Saeed, Zeev Sperber, and Robert C. Valentine. The Intel(R) Pentium(R) M processor: Microarchitecture and performance. Intel Technology Journal, 7(2):21–59, May 2003. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2003/
volume07issue02/art03_pentiumm/p01_abstract.htm.

Silas:2003:SLV

[144] Isic Silas, Igor Frumkin, Eilon Hazan, Ehud Mor, and Genadiy Zobin. System-level validation of the Intel(R) Pentium(R) M processor. Intel Technology Journal, 7(2):37–43, May 2003. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2003/
volume07issue02/art04_validation/p01_abstract.htm.

Genossar:2003:IPM

[145] Dani Genossar and Nachum Shamir. Intel(R) Pentium(R) M processor power estimation, budgeting, optimization and validation. Intel Technology Journal, 7(2):44–49, May 2003. ISSN 1535-766X. URL http://

developer.intel.com/technology/itj/2003/volume07issue02/art05_
power/p01_abstract.htm.

Finkelstein:2003:ASM

[146] Yuval Finkelstein. Antenna selection in multicarrier communication systems. Intel Technology Journal, 7(2):50–58, May 2003. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2003/
volume07issue02/art06_antennas/p01_abstract.htm.

Chao:2003:PWT

[147] Lin Chao. Preface: Wireless technologies. Intel Technology Journal, 7(3): 3, August 2003. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2003/volume07issue03/foreword.htm.

Singer:2003:FWT

[148] Gadi Singer. Foreword: Wireless technologies. Intel Technology Journal, 7(3):4, August 2003. ISSN 1535-766X. URL http://developer.intel.

com/technology/itj/2003/volume07issue03/foreword.htm.

Chesbrough:2003:OPI

[149] Henry Chesbrough. Open Platform Innovation: Creating value from internal and external innovation. Intel Technology Journal, 7(3):5–9, August 2003. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2003/volume07issue03/art01_open/p01_abstract.
htm.

Iyer:2003:PWL

[150] Prakash Iyer, Victor Lortz, Lee Tapper, Roger Chandler, and Roxanne Gryder. Public wireless LAN hotspot deployment and internetworking. Intel Technology Journal, 7(3):10–19, August 2003.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2003/volume07issue03/art02_hotspot/p01_abstract.htm.

Bar-Shalom:2003:UWW

[151] Ofer Bar-Shalom, Gordon Chinn, Kris Fleming, and Uma Gadamsetty. On the union of WPAN and WLAN in mobile computers and handheld devices. Intel Technology Journal, 7(3):20–36, August 2003.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2003/volume07issue03/art03_union/p01_abstract.htm.

Sistla:2003:DWW

[152] Amber Sistla, Jeremy Rover, and Asha Keddy. Dynamic wired and wireless networks on demand. Intel Technology Journal, 7(3): 37–46, August 2003. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2003/volume07issue03/art04_dynamic/
p01_abstract.htm.

Bangerter:2003:HTW

[153] Boyd Bangerter, Eric Jacobsen, Minnie Ho, Adrian Stephens, Alexander Maltsev, Alexey Rubtsov, and Ali Sadri. High-throughput wireless LAN air interface. Intel Technology Journal, 7(3):47–57, August 2003.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2003/volume07issue03/art05_air/p01_abstract.htm.

Chin:2003:SCW

[154] Allan Chin, Ajay Gupta, Ranjit Narjala, and Venkata Vallabhu. Seamless connectivity to wireless local area networks. Intel Technology Journal, 7(3):58–67, August 2003. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2003/volume07issue03/art06_proset/p01_
abstract.htm.

Finnegan:2003:PCP

[155] Jim Finnegan. Preface on communications processing, Intel Technology Journal. Intel Technology Journal, 7(4):3–5, November 2003.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2003/volume07issue04/vol7iss4.pdf.

Maloney:2003:FCP

[156] Sean Maloney. Foreword: Communications processing. Intel Technology Journal, 7(4):6, November 2003. ISSN 1535-766X. URL http://

developer.intel.com/technology/itj/2003/volume07issue04/foreword.
htm.

Boam:2003:MCP

[157] Matthew M. Boam, Jay Gilbert, Tisson K. Mathew, Karel Rasovsky, and Raj Sistla. Modular communications platform. Intel Technology Journal, 7(4):7–16, November 2003. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2003/volume07issue04/art01_modular/
p01_abstract.htm.

Peebles:2003:FAC

[158] Brian Peebles, Chuck Narad, Victoria Genovker, Karel Rasovsky, and Jay Gilbert. Fabrics and application characteristics for AdvancedTCA architectures. Intel Technology Journal, 7(4):17–28, November 2003.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2003/volume07issue04/art02_fabrics/p01_abstract.htm.

Garg:2003:ATC

[159] Sharad Garg, Raj Sistla, Ramesh Caushik, Julie Fleischer, and Rusty Lynch. AdvancedTCA/CGL advantage: HA and TCO. Intel Technology Journal, 7(4):29–38, November 2003. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2003/
volume07issue04/art03_advantage/p01_abstract.htm.

Minturn:2003:ATI

[160] Dave Minturn, Greg Regnier, Jon Krueger, Ravishankar Iyer, and Srihari Makineni. Addressing TCP/IP processing challenges using the IA and IXP processors. Intel Technology Journal, 7(4):39–50, November 2003.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2003/volume07issue04/art04_challenges/p01_abstract.htm.

Deval:2003:DCP

[161] Manasi Deval, Hormuzd Khosravi, Rajeev Muralidhar, Suhail Ahmed, Sanjay Bakshi, and Raj Yavatkar. Distributed control plane architecture for network elements. Intel Technology Journal, 7(4):51–63, November 2003. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2003/volume07issue04/art05_control_plane/p01_
abstract.htm.

Goglin:2003:ASF

[162] Stephen D. Goglin, Donald Hooper, Alok Kumar, and Raj Yavatkar. Advanced software framework, tools, and languages for the IXP family. Intel Technology Journal, 7(4):64–76, November 2003. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2003/
volume07issue04/art06_tools/p01_abstract.htm.

Brink:2003:NPP

[163] Peter Brink, Manohar Castelino, David Meng, Chetan Rawal, and Hari Tadepalli. Network processing performance metrics for the IA- and IXPbased systems. Intel Technology Journal, 7(4):77–91, November 2003.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2003/volume07issue04/art07_metrics/p01_abstract.htm.

Adiletta:2003:EEC

[164] Matthew Adiletta, John Beck, Doug Carrigan, Mark Rosenbluth, Bill Tiso, and Frank Hady. Enterprise edge convergence: Packet processing and computing silicon in the data center. Intel Technology Journal, 7(4):92–106, November 2003. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2003/volume07issue04/art08_convergence/
p01_abstract.htm.

Vinnakota:2003:SII

[165] Bapi Vinnakota, Paul Dormitzer, Mark Rosenbluth, and Sridhar Lakshmanamurthy. Scalable Intel(R) IXA and its building blocks for networking platforms. Intel Technology Journal, 7(4):107–121, November 2003.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2003/volume07issue04/art09_ixa/p01_abstract.htm.

Chao:2004:Pa

[166] Lin Chao. Preface. Intel Technology Journal, 8(1):iii, February 2004. ISSN 1535-766X.

Siu:2004:F

[167] William M. Siu. Foreword. Intel Technology Journal, 8(1):v–vi, February 2004. ISSN 1535-766X.

Boggs:2004:MIP

[168] Darrell Boggs, Aravindh Baktha, Jason Hawkins, Deborah T. Marr, J. Alan Miller, Patrice Roussel, Ronak Singhal, Bret Toll, and K. S. Venkatraman. The microarchitecture of the Intel(R) Pentium(R) 4 processor on 90nm technology. Intel Technology Journal, 8(1):1–17, February 2004. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2004/volume08issue01/foreword.htm.

Smith:2004:SIP

[169] Kevin B. Smith, Aart J. C. Bik, and Xinmin Tian. Support for the Intel(R) Pentium(R) 4 processor with hyper-threading technology in Intel(R) 8.0 compilers. Intel Technology Journal, 8(1):19–31, February 2004.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2004/volume08issue01/art02_compilers/p01_abstract.htm.

Singhal:2004:PAV

[170] Ronak Singhal, K. S. Venkatraman, . Evan R. Cohn, John G. Holm, David A. Koufaty, Meng-Jang Lin, Mahesh J. Madhav, Markus Mattwandel, Nidhi Nidhi, Jonathan D. Pearce, and Madhusudanan Seshadri. Performance analysis and validation of the Intel(R) Pentium(R) 4 processor on 90nm technology. Intel Technology Journal, 8(1):33–42, February 2004.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2004/volume08issue01/art03_performance/p01_abstract.htm.

Deleganes:2004:LTI

[171] Dan J. Deleganes, Micah Barany, Daniel Chow, Tom D. Fletcher, George L. Geannopoulos, Kurt Kreitzer, Anant P. Singh, and Sapumal B. Wijeratne. LVS technology for the Intel(R) Pentium(R) 4 processor on 90nm technology. Intel Technology Journal, 8(1):43–53, February 2004.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2004/volume08issue01/art04_lvs_technology/p01_abstract.htm.

Chappell:2004:LAC

[172] Barbara Chappell, Amanda Duncan, Kiran Ganesh, Manoj Gunwani, Abhinav Sharma, and Desktop Madhu Swarna. Library architecture challenges for cell-based design. Intel Technology Journal, 8(1): 55–61, February 2004. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2004/volume08issue01/art05_library_
arch/p01_abstract.htm.

Kuppuswamy:2004:FHS

[173] Ravishankar Kuppuswamy, Peter DesRosier, Derek Feltham, Rehan Sheikh, and Paul Thadikaran. Full hold-scan systems in microprocessors: Cost/benefit analysis. Intel Technology Journal, 8(1):63–, February 2004.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2004/volume08issue01/art06_full_hold/p01_abstract.htm.

Chao:2004:Pb

[174] Lin Chao. Preface. Intel Technology Journal, 8(2):iii–iv, May 2004. ISSN 1535-766X.

Paniccia:2004:NEO

[175] Mario Paniccia. Foreword: a new era in optical communications. Intel Technology Journal, 8(2):v–vi, May 2004. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2004/
volume08issue02/foreword.htm.

Anonymous:2004:TRa

[176] Anonymous. Technical reviewers. Intel Technology Journal, 8(2):vii, May 2004. ISSN 1535-766X.

Herve:2004:GOT

[177] Pierre Herve and Shlomo Ovadia. Optical technologies for enterprise networks. Intel Technology Journal, 8(2):73–82, May 2004. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2004/
volume08issue02/art02_opttransceivers/p01_abstract.htm.

Kirkpatrick:2004:GOT

[178] Peter Kirkpatrick, Wei chiao Fang, Henrik Johansen, Benny Christensen, Jesper Hanberg, Martin Lobel, Tom Mader, Song Shang, Craig Schulz, Doug Sprock, and Marc Verdiell. 10 Gb/s optical transceivers: Fundamentals and emerging technologies. Intel Technology Journal, 8(2): 83–99, May 2004. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2004/volume08issue02/art02_opttransceivers/p01_
abstract.htm.

Finot:2004:AOP

[179] Marc Finot, Mark McDonald, Andrew Daiber, William B. Chapman, Delin Li, Marc Epitaux, Eric Zbinden, Jeff Bennett, William J. Kozlovsky, and Jean-Marc Verdiell. Automated optical packaging technology for 10 Gb/s transceivers and its application to a low-cost full C-band tunable transmitter. Intel Technology Journal, 8(2):101–114,

May 2004. ISSN 1535-766X. URL http://developer.intel.com/
technology/itj/2004/volume08issue02/art03_packagingtech/p01_
abstract.htm.

Mohammed:2004:OIS

[180] Edris Mohammed, Andrew Alduino, Thomas Thomas, Henning Braunisch, Daoqiang Lu, John Heck, Ansheng Liu, Ian Young, Brandon Barnett, Gilroy Vandentop, and Randy Mooney. Optical interconnect system integration for ultra-short-reach applications. Intel Technology Journal, 8(2):115–127, May 2004. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2004/volume08issue02/art04_interconnect/
p01_abstract.htm.

Kobrinsky:2004:COI

[181] Mauro J. Kobrinsky, Bruce A. Block, Jun-Fei Zheng, Brandon C. Barnett, Edris Mohammed, Miriam Reshotko, Frank Robertson, Scott List, Ian Young, and Kenneth Cadien. On-chip optical interconnects. Intel Technology Journal, 8(2):129–141, May 2004. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2004/
volume08issue02/art05_on-chip/p01_abstract.htm.

Kobrinsky:2004:SP

[182] Mike Salib, Ling Liao, Richard Jones, Mike Morse, Ansheng Liu, Dean Samara-Rubio, Drew Alduino, and Mario Paniccia. Silicon photonics. Intel Technology Journal, 8(2):143–160, May 2004. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2004/
volume08issue02/art06_siliconphoto/p01_abstract.htm.

Hutchinson:2004:IPB

[183] John M. Hutchinson, Jun-Fei Zheng, Jonathon S. Barton, Jeffrey A. Henness, Milan L. Maaanovi, Matthew N. Sysak, Leif A. Johansson, Daniel J. Blumenthal, Larry A. Coldren, Hilmi Volkan Demir, Edward. L. Ginzton, Vijit A. Sabnis, Onur Fidaner, James S. Harris, and David A. B. Miller. Indium phosphide-based optoelectronic wavelength conversion for highspeed optical networks. Intel Technology Journal, 8(2):161–171, May 2004.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2004/volume08issue02/art07_indium/p01_abstract.htm.

Chao:2004:Pc

[184] Lin Chao. Preface. Intel Technology Journal, 8(3):iii–iv, August 2004. ISSN 1535-766X.

Richardson:2004:FEB

[185] Scott G. Richardson. Foreword: Emerging broadband networks: The case for WiMAX. Intel Technology Journal, 8(3):v–vi, August 2004.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2004/volume08issue03/foreword.htm.

Anonymous:2004:TRb

[186] Anonymous. Technical reviewers. Intel Technology Journal, 8(3):vii, August 2004. ISSN 1535-766X.

Agis:2004:GIB

[187] Ed Agis, Henry Mitchel, Shlomo Ovadia, Selim Aissi, Sanjay Bakshi, Prakash Iyer, Masud Kibria, Christopher Rogers, and James Tsai. Global, interoperable broadband wireless networks: Extending WiMAX technology to mobility. Intel Technology Journal, 8(3):173–187, Au-

gust 2004. ISSN 1535-766X. URL http://developer.intel.com/
technology/itj/2004/volume08issue03/art01_globalwirelessnet/
p01_abstract.htm.

Bisla:2004:RSC

[188] Balvinder Bisla, Roger Eline, and Luiz M. Franca-Neto. RF system and circuit challenges for WiMAX. Intel Technology Journal, 8(3):189–200, August 2004. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2004/volume08issue03/art02_rfsystem/p01_abstract.
htm.

Yaghoobi:2004:SOP

[189] Hassan Yaghoobi. Scalable OFDMA physical layer in IEEE 802.16 WirelessMAN. Intel Technology Journal, 8(3):201–212, August 2004. ISSN

1535-766X.  URL http://developer.intel.com/technology/itj/
2004/volume08issue03/art03_scalableofdma/p01_abstract.htm.

Nair:2004:IMA

[190] Govindan Nair, Joey Chou, Tomasz Madejski, Krzysztof Perycz, David Putzolu, and Jerry Sydir. IEEE 802.16 medium access control and service provisioning. Intel Technology Journal, 8(3):213–228, August 2004.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2004/volume08issue03/art04_ieee80216mac/p01_abstract.htm.

Salvekar:2004:MAT

[191] Atul Salvekar, Sumeet Sandhu, Qinghua Li, Minh-Anh Vuong, and Xiaoshu Qian. Multiple antenna technology in WiMAX systems. Intel Technology Journal, 8(3):229–239, August 2004. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2004/
volume08issue03/art05_multiantenna/p01_abstract.htm.

Franca-Neto:2004:FIC

[192] Luiz M. Franca-Neto, Roger Eline, and Bisla Balvinder. Fully integrated CMOS radios from RF to millimeter wave frequencies. Intel Technology Journal, 8(3):241–258, August 2004. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2004/
volume08issue03/art06_intcmosradios/p01_abstract.htm.

Chao:2004:Pd

[193] Lin Chao. Preface. Intel Technology Journal, 8(4):iii, November 2004.

ISSN 1535-766X.

Busch:2004:FBI

[194] Doug Busch. Foreword: Building Intel leadership in enterprise computing. Intel Technology Journal, 8(4):v–vi, November 2004. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2004/
volume08issue04/foreword.htm.

Anonymous:2004:TRc

[195] Anonymous. Technical reviewers. Intel Technology Journal, 8(4):vii, November 2004. ISSN 1535-766X.

Brown:2004:PE

[196] George Brown, Thomas Gardos, Jay Hopman, Hong Li, Sigal Louchheim, Cynthia Pickering, Jeff Sedayao, and John Vicente. The proactive enterprise. Intel Technology Journal, 8(4):259–267, November 2004.

ISSN 1535-766X. URL ftp://download.intel.com/technology/itj/
2004/volume08issue04/art01_proactive/vol8_art01.pdf;   http:
//developer.intel.com/technology/itj/2004/volume08issue04/art01_
proactive/p01_abstract.htm.

Sedayao:2004:PAP

[197] Jeff Sedayao, John Vicente, Rita Wouhaybi, Hong Li, Manish Dave, Sanjay Rungta, and Stacy Purcell. PlanetLab and its applicability to the Proactive Enterprise. Intel Technology Journal, 8(4):269–277, November 2004. ISSN 1535-766X. URL http://developer.intel.

com/technology/itj/2004/volume08issue04/art02_planetlab/p01_
abstract.htm.

Melcher:2004:TAF

[198] Brian Melcher and Bradley Mitchell. Towards an autonomic framework: Self-configuring network services and developing autonomic applications. Intel Technology Journal, 8(4):279–290, November 2004.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2004/volume08issue04/art03_autonomic/p01_abstract.htm.

Krishnaswamy:2004:SAW

[199] Dilip Krishnaswamy and John Vicente. Scalable adaptive wireless networks for multimedia in the proactive enterprise. Intel Technology Journal, 8(4):291–301, November 2004. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2004/
volume08issue04/art04_scalingwireless/p01_abstract.htm.

Rungta:2004:BSP

[200] Sanjay Rungta, Anant Raman, Toby Kohlenberg, Hong Li, Manish Dave, and Greg Kime. Bringing security proactively into the enterprise. Intel Technology Journal, 8(4):303–311, November 2004.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2004/volume08issue04/art05_security/p01_abstract.htm.

Kozuch:2004:ECM

[201] Michael A. Kozuch, Casey J. Helfrich, David O. Hallaron, and Mahadev Satyanarayanan. Enterprise client management with Internet Suspend/Resume. Intel Technology Journal, 8(4):313–323, Novem-

ber 2004. ISSN 1535-766X. URL http://developer.intel.com/
technology/itj/2004/volume08issue04/art06_suspendresume/p01_
abstract.htm.

Goodwin:2004:AAS

[202] Randall Goodwin, Russell Miller, Eugene Tuv, Alexander Borisov, Mani Janakiram, and Sigal Louchheim. Advancements and applications of statistical learning/data mining in semiconductor manufacturing. Intel Technology Journal, 8(4):325–336, November 2004.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2004/volume08issue04/art07_datamining/p01_abstract.htm.

Thangarathinam:2004:MMF

[203] Thiru Thangarathinam, Gregg Wyant, Jacque Gibson, and John Simpson. Metadata management: the foundation for enterprise information integration. Intel Technology Journal, 8(4):337–344, November 2004.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2004/volume08issue04/art08_metadata/p01_abstract.htm.

Brown:2004:SAS

[204] George Brown and Robert Carpenter. Successful application of service-oriented architecture across the enterprise and beyond. Intel Technology Journal, 8(4):345–359, November 2004. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2004/
volume08issue04/art09_successful/p01_abstract.htm.

Agosta:2004:BNS

[205] John Mark Agosta and Thomas Gardos. Bayes network “smart” diagnostics. Intel Technology Journal, 8(4):361–372, November 2004.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2004/volume08issue04/art10_bayesnetwork/p01_abstract.htm.

Pickering:2004:ABP

[206] Cynthia Pickering and Eleanor Wynn. An architecture and business process framework for global team collaboration. Intel Technology Journal, 8 (4):373–382, November 2004. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2004/volume08issue04/art11_collaboration/
p01_abstract.htm.

Chao:2005:PSG

[207] Lin Chao. Preface: Second-generation Intel(R) Centrino^{T M} mobile technology. Intel Technology Journal, 9(1):iii–iv, February 17, 2005. ISSN 1535-766X.

Anonymous:2005:TRa

[208] Anonymous. Technical reviewers. Intel Technology Journal, 9(1):v, February 17, 2005. ISSN 1535-766X.

Thakkar:2005:SGI

[209] Shreekant (Ticky) Thakkar. Second-generation Intel(R) Centrino^{T M} mobile technology platform. Intel Technology Journal, 9(1):1–10, February 17, 2005. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2005/volume09issue01/intro/p01_abstract.htm.

Browning:2005:ICM

[210] David W. Browning, Vinay Kumar C., Praveen G., Ananth V. Gopal, Tom Shewchuk, and Suresh V. Subramanyam. Intel(R) 915GMS chipset: In mobile platforms, smaller is better. Intel Technology Journal, 9(1):11–19, February 17, 2005. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2005/
volume09issue01/art01_915gms_chipset/p01_abstract.htm.

Kolbehdari:2005:EPE

[211] Mohammad Kolbehdari, David Harriman, Altug Koker, Seh Kwa, and Brad Saunders. The emergence of PCI Express* in the next generation of mobile platforms. Intel Technology Journal, 9(1):21–33, February 17, 2005.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2005/volume09issue01/art02_pcix_mobile/p01_abstract.htm.

Leung:2005:HPG

[212] Kam Leung, Steve Spangler, Todd Witter, Kevin E. Arendt, Satya Avadhanam, Satyaki Koneru, and Val Cook. High-performance graphics and TV output comes to the second-generation Intel(R) Centrino^{T M} mobile technology platform. Intel Technology Journal, 9(1):35–47, February 17, 2005. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2005/volume09issue01/art03_hiperf_graphics/p01_
abstract.htm.

Clark:2005:LPA

[213] Tom Clark, Feng (Frank) Yang, Greg Kaine, Brian Leete, and Sriram Ranganathan. Low-power audio and storage input/output technologies for the second-generation Intel(R) Centrino^{T M} mobile technology platform. Intel Technology Journal, 9(1):49–59, February 17, 2005.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2005/volume09issue01/art04_audio_iotech/p01_abstract.htm.

Chang:2005:PPC

[214] Eunice Chang, Jim Fung, Greg Kaine, Ali Saeed, and Randy P. Stanley. Performance and power consumption for mobile platform components under common usage models. Intel Technology Journal, 9(1):61–73, February 17, 2005. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2005/volume09issue01/art05_perf_power/p01_abstract.
htm.

Samson:2005:IMS

[215] Eric C. Samson, Sridhar V. Machiroutu, Je-Young Chang, Ishmael Santos, Jim Hermerding, Ashay Dani, Ravi Prasher, and David W. Song. Interface material selection and a thermal management technique in secondgeneration platforms built on Intel(R) Centrino^{T M} mobile technology. Intel Technology Journal, 9(1):75–86, February 17, 2005. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2005/
volume09issue01/art06_interface_materials/p01_abstract.htm.

Chen:2005:NGP

[216] Camille Chen, Penny Chen, Sherry Chen, Gordon Chinn, Hani Elgebaly, Uma Gadamsetty, Ranjit Narjala, Paul K. Reddy, Jeremy Rover, and Paul Schmitz. Next-generation PC platform built on Intel(R) Centrino^{T M} mobile technology new usage models. Intel Technology Journal, 9(1):87–98, February 17, 2005. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2005/
volume09issue01/art07_next_generation/p01_abstract.htm.

Chao:2005:PCI

[217] Lin Chao. Preface: Compute-intensive, highly parallel applications and uses. Intel Technology Journal, 9(2):iii–iv, May 2005. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2005/
volume09issue02/preface.htm.

Liang:2005:FRM

[218] Bob Liang and Pradeep Dubey. Foreword: Recognition, mining and synthesis. Intel Technology Journal, 9(2):v–vi, May 2005. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2005/
volume09issue02/foreword.htm.

Hurley:2005:RTG

[219] Jim Hurley. Ray tracing goes mainstream. Intel Technology Journal, 9(2): 99–107, May 2005. ISSN 1535-766X. URL ftp://download.intel.com/

technology/itj/2005/volume09issue02/art01_ray_tracing/vol09_
art01.pdf;   http://developer.intel.com/technology/itj/2005/
volume09issue02/art01_ray_tracing/p01_abstract.htm.

Chen:2005:CVW

[220] Trista P. Chen, Horst Haussecker, Alexander Bovyrin, Roman Belenov, Konstantin Rodyushkin, Alexander Kuranov, and Victor Eruhimov. Computer vision workload analysis: Case study of video surveillance systems. Intel Technology Journal, 9(2):109–118, May 2005. ISSN

1535-766X. URL ftp://download.intel.com/technology/itj/2005/
volume09issue02/art02_computer_vision/vol09_art02.pdf;   http:
//developer.intel.com/technology/itj/2005/volume09issue02/art02_
computer_vision/p01_abstract.htm.

Bradski:2005:LBC

[221] Gary Bradski, Adrian Kaehler, and Vadim Pisarevsky. Learningbased computer vision with Intel’s open source computer vision library. Intel Technology Journal, 9(2):119–130, May 2005. ISSN

1535-766X. URL ftp://download.intel.com/technology/itj/2005/
volume09issue02/art03_learning_vision/vol09_art03.pdf;   http:
//developer.intel.com/technology/itj/2005/volume09issue02/art03_
learning_vision/p01_abstract.htm.

Chen:2005:PSD

[222] Yurong Chen, Qian Diao, Carole Dulong, Chunrong Lai, Wei Hu, Eric Li, Wenlong Li, Tao Wang, and Yimin Zhang. Performance scalability of data-mining workloads in bioinformatics. Intel Technology Journal, 9(2): 131–142, May 2005. ISSN 1535-766X. URL ftp://download.intel.com/

technology/itj/2005/volume09issue02/art04_data_workloads/vol09_
art04.pdf;   http://developer.intel.com/technology/itj/2005/
volume09issue02/art04_data_workloads/p01_abstract.htm.

Borisov:2005:PSA

[223] Alexander Borisov, Igor Chikalov, Victor Eruhimov, and Eugene Tuv. Performance and scalability analysis of tree-based models in largescale data-mining problems. Intel Technology Journal, 9(2):143–150,

May 2005. ISSN 1535-766X. URL ftp://download.intel.com/
technology/itj/2005/volume09issue02/art05_tree-based_models/
vol09_art05.pdf;   http://developer.intel.com/technology/itj/
2005/volume09issue02/art05_tree-based_models/p01_abstract.htm.

Smelyanskiy:2005:PCL

[224] Mikhail Smelyanskiy, Stephen Skedzielewski, and Carole Dulong. Parallel computing for large-scale optimization problems: Challenges and solutions. Intel Technology Journal, 9(2):151–163, May 2005. ISSN

1535-766X. URL ftp://download.intel.com/technology/itj/2005/
volume09issue02/art06_parallel_computing/vol09_art06.pdf;   http:
//developer.intel.com/technology/itj/2005/volume09issue02/art06_
parallel_computing/p01_abstract.htm.

Srinivasan:2005:UPR

[225] Krishnamurthy Srinivasan, Raj Ramanujan, Michael Amirfathi, and Enrique Castro-Leon. Understanding the platform requirements of emerging enterprise solutions. Intel Technology Journal, 9(2):165–175, May

2005. ISSN 1535-766X. URL ftp://download.intel.com/technology/
itj/2005/volume09issue02/art07_platform_req/vol09_art07.pdf;
 http://developer.intel.com/technology/itj/2005/volume09issue02/
art07_platform_req/p01_abstract.htm.

Chao:2005:PMI

[226] Lin Chao. Preface: Managing international supply and demand at Intel Corporation. Intel Technology Journal, 9(3):iii–v, August 2005.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2005/volume09issue03/preface.htm.

Kempf:2005:FMI

[227] Karl Kempf. Foreword: Managing Intel’s international network of supply and demand. Intel Technology Journal, 9(3):vii–viii, August 2005.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2005/volume09issue03/foreword.htm.

Anonymous:2005:TRb

[228] Anonymous. Technical reviewers. Intel Technology Journal, 9(3):ix, August 2005. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2005/volume09issue03/foreword.htm.

Hopman:2005:MUP

[229] Jay W. Hopman. Managing uncertainty in planning and forecasting. Intel Technology Journal, 9(3):175–183, August 2005. ISSN

1535-766X. URL ftp://download.intel.com/technology/itj/2005/
volume09issue03/art01_mnguncertainty/vol09_art01.pdf;   http:
//developer.intel.com/technology/itj/2005/volume09issue03/art01_
mnguncertainty/p01_abstract.htm.

Vaidyanathan:2005:UCO

[230] Viswanath Vaidyanathan, Dave Metcalf, and Douglas Martin. Using capacity options to better enable our factory ramps. Intel Technology Journal, 9(3):185–191, August 2005. ISSN 1535-766X. URL ftp://download.

intel.com/technology/itj/2005/volume09issue03/art02_capacityopts/
vol09_art02.pdf;   http://developer.intel.com/technology/itj/
2005/volume09issue03/art02_capacityopts/p01_abstract.htm.

Katz:2005:RTE

[231] Jason Katz and Don Edenfeld. Redefining the test equipment supply chain: The open architecture revolution. Intel Technology Journal, 9(3): 193–201, August 2005. ISSN 1535-766X. URL ftp://download.intel.

com/technology/itj/2005/volume09issue03/art03_testequipment/
vol09_art03.pdf;   http://developer.intel.com/technology/itj/
2005/volume09issue03/art03_testequipment/p01_abstract.htm.

Ghiya:2005:PSI

[232] Kinnar Ghiya and Marci Powers. e-procurement — strengthening the indirect supply-chain through technology globalization. Intel Technology Journal, 9(3):203–209, August 2005. ISSN 1535-766X. URL ftp://download.

intel.com/technology/itj/2005/volume09issue03/art04_eprocurement/
vol09_art04.pdf;   http://developer.intel.com/technology/itj/
2005/volume09issue03/art04_eprocurement/p01_abstract.htm.

Ali:2005:IPC

[233] Anwar Ali, William J. Campbell, Lance I. Solomon, Megan A. Walsh, and James R. Wuerfel. Intel’s processes for capacity planning optimization. Intel Technology Journal, 9(3):211–221, August 2005. ISSN

1535-766X. URL ftp://download.intel.com/technology/itj/2005/
volume09issue03/art05_intelsprocess/vol09_art05.pdf;   http:
//developer.intel.com/technology/itj/2005/volume09issue03/art05_
intelsprocess/p01_abstract.htm.

Bean:2005:OSC

[234] John W. Bean, Amit Devpura, Michael O. Brien, and Shamin Shirodkar. Optimizing supply-chain planning. Intel Technology Journal, 9(3):223– 231, August 2005. ISSN 1535-766X. URL ftp://download.intel.com/

technology/itj/2005/volume09issue03/art06_supplychain/vol09_
art06.pdf;   http://developer.intel.com/technology/itj/2005/
volume09issue03/art06_supplychain/p01_abstract.htm.

Johnson:2005:IM

[235] Kurt L. Johnson. Inventory modeling. Intel Technology Journal, 9(3):233– 238, August 2005. ISSN 1535-766X. URL ftp://download.intel.com/

technology/itj/2005/volume09issue03/art07_inventorymodeling/
vol09_art07.pdf;   http://developer.intel.com/technology/itj/
2005/volume09issue03/art07_inventorymodeling/p01_abstract.htm.

Cartwright:2005:RIT

[236] John Cartwright, Jay Hahn-Steichen, Jackson He, and Thurman Miller. RosettaNet for Intel’s trading entity automation. Intel Technology Journal, 9(3):239–245, August 2005. ISSN 1535-766X. URL ftp://download.

intel.com/technology/itj/2005/volume09issue03/art08_rosettanet/
vol09_art08.pdf;   http://developer.intel.com/technology/itj/
2005/volume09issue03/art08_rosettanet/p01_abstract.htm.

Dighero:2005:RRI

[237] Craig Dighero, James Kellso, Debbie Merizon, Mary Murphy-Hoye, and Richard Tyo. RFID: The real and integrated story. Intel Technology Journal, 9(3):247–257, August 2005. ISSN 1535-766X. URL ftp://download.

intel.com/technology/itj/2005/volume09issue03/art09_rfid/vol09_
art09.pdf;   http://developer.intel.com/technology/itj/2005/
volume09issue03/art09_rfid/p01_abstract.htm.

Chao:2005:PEP

[238] Lin Chao. Preface: Electronic package technology development. Intel Technology Journal, 9(4):iii, November 9, 2005. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2005/
volume09issue04/preface.htm.

Grayeli:2005:FCO

[239] Nasser Grayeli. Foreword: Challenges and opportunities in electronic package technology development. Intel Technology Journal, 9(4):v–vi, November 9, 2005. ISSN 1535-766X. URL http://developer.intel.

com/technology/itj/2005/volume09issue04/foreword.htm.

Anonymous:2005:TRc

[240] Anonymous. Technical reviewers. Intel Technology Journal, 9(4):vii, November 9, 2005. ISSN 1535-766X. URL http://developer.intel.

com/technology/itj/2005/volume09issue04/foreword.htm.

Mallik:2005:APT

[241] Debendra Mallik, Kaladhar Radhakrishnan, Jinagqi He, Chia-Pin Chiu, Telesphor Kamgaing, Damion Searls, and James D. Jackson. Advanced package technologies for high-performance systems. Intel Technology Journal, 9(4):259–271, November 9, 2005. ISSN

1535-766X. URL ftp://download.intel.com/technology/itj/2005/
volume09issue04/art01_advpackagetech/vol09_art01.pdf;   http:
//developer.intel.com/technology/itj/2005/volume09issue04/art01_
advpackagetech/p01_abstract.htm.

Aygun:2005:PDH

[242] Kemal Aygün, Michael J. Hill, Kimberley Eilert, Kaladhar Radhakrishnan, and Alex levin. Power delivery for high-performance microprocessors. Intel Technology Journal, 9(4):273–283, November 9, 2005. ISSN

1535-766X. URL ftp://download.intel.com/technology/itj/2005/
volume09issue04/art02_powerdelivery/vol09_art02.pdf;   http:
//developer.intel.com/technology/itj/2005/volume09issue04/art02_
powerdelivery/p01_abstract.htm.

Prasher:2005:NMT

[243] Ravi S. Prasher, Je-Young Chang, Ioan Sauciuc, Sridhar Narasimhan, David Chau, Greg Chrysler, Alan Myers, Suzana Prstic, and Chuan Hu. Nano and micro technology-based next-generation package-level cooling solutions. Intel Technology Journal, 9(4):285–296, November 9,

2005. ISSN 1535-766X. URL ftp://download.intel.com/technology/
itj/2005/volume09issue04/art03_nanoandmicro/vol09_art03.pdf;
 http://developer.intel.com/technology/itj/2005/volume09issue04/
art03_nanoandmicro/p01_abstract.htm.

Garner:2005:FSC

[244] Luke Garner, Sandeep Sane, Daewoong Suh, Tiffany Byrne, Ashay Dani, Ted Martin, Michael Mello, Mitesh Patel, and Richard Williams. Finding solutions to the challenges in package interconnect reliability. Intel Technology Journal, 9(4):297–308, November 9, 2005.

ISSN 1535-766X. URL ftp://download.intel.com/technology/itj/
2005/volume09issue04/art04_findingsolutions/vol09_art04.pdf;
 http://developer.intel.com/technology/itj/2005/volume09issue04/
art04_findingsolutions/p01_abstract.htm.

Wakharkar:2005:MTT

[245] Vijay Wakharkar, Chris Matayabas, Ed Lehman, Rahul Manepalli, Mukul Renavikar, Saikumar Jayaraman, and Vassou LeBonheur. Materials technologies for thermomechanical management of organic packages. Intel Technology Journal, 9(4):309–323, November 9, 2005. ISSN

1535-766X. URL ftp://download.intel.com/technology/itj/2005/
volume09issue04/art05_materialstech/vol09_art05.pdf;   http:

//developer.intel.com/technology/itj/2005/volume09issue04/art05_
materialstech/p01_abstract.htm.

Baldwin:2005:PPH

[246] Chris Baldwin, Tod Byquist, Chris Combs, Vinayak Pandey, Brent Stone, Ram Viswanath, Abhay Watwe, and Leight Wojewoda. Pentium 4 processor high-volume land-grid-array technology: Challenges and future trends. Intel Technology Journal, 9(4):325–336, November 9, 2005. ISSN

1535-766X. URL ftp://download.intel.com/technology/itj/2005/
volume09issue04/art06_pentium4proc/vol09_art06.pdf;   http:/
/developer.intel.com/technology/itj/2005/volume09issue04/art06_
pentium4proc/p01_abstract.htm.

Pacheco:2005:AFI

[247] Mario Pacheco, Zhiyong Wang, Lars Skoglund, Yongmei Liu, Ariel Medina, Arun Raman, Rajen Dias, Deepak Goyal, and Shriram Ramanathan. Advanced fault isolation and failure analysis techniques for future package technologies. Intel Technology Journal, 9(4):337–352, November 9, 2005. ISSN 1535-766X. URL ftp://download.intel.com/

technology/itj/2005/volume09issue04/art07_advfaultisolation/
vol09_art07.pdf;   http://developer.intel.com/technology/itj/
2005/volume09issue04/art07_advfaultisolation/p01_abstract.htm.

Kamgaing:2005:FPT

[248] Telesphor Kamgaing, Kinya Ichikawa, Xiang Yin Zeng, Kyu-Pyung Hwang, Yonkgki Min, and Jiro Kubota. Future package technologies for wireless communication systems. Intel Technology Journal, 9(4):353–364, November 9, 2005. ISSN 1535-766X. URL ftp://download.intel.com/

technology/itj/2005/volume09issue04/art08_futurepackagetech/
vol09_art08.pdf;   http://developer.intel.com/technology/itj/
2005/volume09issue04/art08_futurepackagetech/p01_abstract.htm.

Chao:2006:PVV

[249] Lin Chao. Preface: Voice, video, and data services — the merging of triplets. Intel Technology Journal, 10(1):iii–iv, February 2006.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2006/volume10issue01/preface.htm.

Neal-Graves:2006:FCN

[250] Anthony Neal-Graves. Foreword: Convergence now. Intel Technology Journal, 10(1):v–vi, February 2006. ISSN 1535-766X. URL http://

developer.intel.com/technology/itj/2006/volume10issue01/foreword.
htm.

Anonymous:2006:TRa

[251] Anonymous. Technical reviewers. Intel Technology Journal, 10(1):v, February 2006. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2006/volume10issue01/foreword.htm.

Rungta:2006:ECN

[252] Sanjay Rungta and Omer Ben-Shalom. Enterprise converged network-one network for voice, video, data, and wireless. Intel Technology Journal, 10(1):1–9, February 2006. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2006/volume10issue01/art01_enterprise_
converged_network/p01_abstract.htm;   http://www.intel.com/technology/
itj/2006/volume10issue01/art01_enterprise_converged_network/
vol10_art01.pdf.

Kolbehdari:2006:SIP

[253] Mohammad Kolbehdari, Dave Lizotte, Glen Shires, and Scott Trevor. Session Initiation Protocol (SIP) evolution in converged communications. Intel Technology Journal, 10(1):11–18, February 2006. ISSN

1535-766X.  URL http://developer.intel.com/technology/itj/
2006/volume10issue01/art02_sip_evolution/p01_abstract.htm;
http://www.intel.com/technology/itj/2006/volume10issue01/art02_
sip_evolution/vol10_art02.pdf.

DeNies:2006:SBI

[254] Steve DeNies, Jay Gilbert, Ashok Mishra, and Karel Rasovsky. Standardsbased interoperability for the Advanced Telecom Computing Architecture (AdvancedTCA*). Intel Technology Journal, 10(1):19–27, February 2006.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2006/volume10issue01/art03_advanced_tca_interoperability/
p01_abstract.htm;   http://www.intel.com/technology/itj/2006/
volume10issue01/art03_advanced_tca_interoperability/vol10_art03.
pdf.

Sinha:2006:QCV

[255] Ranjan Sinha, Catherine Spence, and Tim Verrall. Quality campus VoIP: An Intel case study. Intel Technology Journal, 10(1): 29–38, February 2006. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2006/volume10issue01/art04_quality_
campus_voip/p01_abstract.htm;   http://www.intel.com/technology/
itj/2006/volume10issue01/art04_quality_campus_voip/vol10_art04.
pdf.

Siegel:2006:EPB

[256] Shelby Siegel, Dave Lizotte, Blaine Bauer, Maria Frick, and Duncan Glendinning. Experiences with PC-based real-time multimedia collaboration over IP. Intel Technology Journal, 10(1):39–47, February 2006. ISSN

1535-766X.  URL http://developer.intel.com/technology/itj/
2006/volume10issue01/art05_real_time_multimedia_collaboration/
p01_abstract.htm;   http://www.intel.com/technology/itj/2006/
volume10issue01/art05_real_time_multimedia_collaboration/vol10_
art05.pdf.

Grecco:2006:UIT

[257] Joseph Grecco, Mark Mize, and Ranjan Singh. Using Intel technologies to build next-generation media servers. Intel Technology Journal, 10(1): 49–66, February 2006. ISSN 1535-766X. URL http://developer.intel.

com/technology/itj/2006/volume10issue01/art06_next_generation_
media_servers/p01_abstract.htm;   http://www.intel.com/technology/
itj/2006/volume10issue01/art06_next_generation_media_servers/
vol10_art06.pdf.

Walker:2006:NUP

[258] Mark Walker, Jim Edwards, Ylian Saint-Hilaire, and Jeffrey Lee. New uses, proposed standards, and emergent device classes for digital home communications. Intel Technology Journal, 10(1):67–75, February 2006.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2006/volume10issue01/art07_uses_standards_device_classes/
p01_abstract.htm;   http://www.intel.com/technology/itj/2006/
volume10issue01/art07_uses_standards_device_classes/vol10_art07.
pdf.

Elgebaly:2006:SCE

[259] Hani Elgebaly, Farid Adrangi, Rajeev Muralidhar, Lakshmi Ramachandran, and Manish Nair. Seamless collaboration-enabling best-in-class VoIP experience on Intel Centrino mobile technology. Intel Technology Journal, 10(1):77–87, February 2006. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2006/
volume10issue01/art08_seamless_voip_centrino_mobile/p01_abstract.
htm;   http://www.intel.com/technology/itj/2006/volume10issue01/
art08_seamless_voip_centrino_mobile/vol10_art08.pdf.

Chao:2006:PMI

[260] Lin Chao. Preface: Meet the Intel Core Duo processor. Intel Technology Journal, 10(2):iii–iv, May 15, 2006. ISSN 1535-766X. URL http://

developer.intel.com/technology/itj/index.htm.

Perlmutter:2006:FIC

[261] Dadi Perlmutter. Foreword: Intel Centrino Duo mobile technology: The beginning of an era of mobile multi-core computing. Intel Technology Journal, 10(2):v–vi, May 15, 2006. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2006/volume10issue02/foreword.htm.

Anonymous:2006:TRb

[262] Anonymous. Technical reviewers. Intel Technology Journal, 10(2):vii, May 15, 2006. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2006/volume10issue02/foreword.htm.

Gochman:2006:IIC

[263] Simcha Gochman, Avi Mendelson, Alon Naveh, and Efraim Rotem. Introduction to Intel Core Duo processor architecture. Intel Technology Journal, 10(2):89–97, May 15, 2006. ISSN 1535-766X. URL http://

developer.intel.com/technology/itj/2006/volume10issue02/art01_
Intro_to_Core_Duo/p01_abstract.htm.

Mendelson:2006:CIS

[264] Avi Mendelson, Julius Mandelblat, Simcha Gochman, Anat Shemer, Erik Niemeyer, and Arun Kumar. CMP implementation in systems based on the Intel Core Duo processor. Intel Technology Journal, 10(2):99–107, May 15, 2006. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2006/volume10issue02/art02_CMP_Implementation/
p01_abstract.htm.

Naveh:2006:PTM

[265] Alon Naveh, Efraim Rotem, Avi Mendelson, Simcha Gochman, Rajshree Chabukswar, Karthik Krishnan, and Arun Kumar. Power and thermal management in the Intel Core Duo processor. Intel Technology Journal, 10(2):109–122, May 15, 2006. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2006/
volume10issue02/art03_Power_and_Thermal_Management/p01_abstract.
htm.

Iyer:2006:SMP

[266] Jayesh Iyer, Corinne L. Hall, Jerry Shi, and Yuchen Huang. System memory power and thermal management in platforms built on Intel Centrino Duo mobile technology. Intel Technology Journal, 10(2):123–132, May 15, 2006. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2006/volume10issue02/art04_Memory_Power_Management/
p01_abstract.htm.

Subramanyam:2006:IEC

[267] Suresh Subramanyam, Taninder Sijher, Sidharth Krishnama, Parthasarathy Ramaswamy, Deepa Mohan, Vikas Shilimkar, Eric Samson, Michael Derr, and Samir Gundawar. Intel 945GMS Express chipset for small form factor platform based on Intel Centrino Duo mobile technology. Intel Technology Journal, 10(2):133–145, May 15, 2006. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2006/
volume10issue02/art05_945GMS_SFF_Low_Voltage/p01_abstract.htm.

Ruberto:2006:WSH

[268] Mark Ruberto, Ra’anan Soyer, Jorge Myszne, Alexander Sloutsky, and Yari Shemesh. WLAN system, HW, and RFIC architecture for the Intel PRO/Wireless 3945ABG network connection. Intel Technology Journal, 10(2):147–156, May 15, 2006. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2006/volume10issue02/art06_WLAN_HW_
RFIC/p01_abstract.htm.

Hen:2006:MAW

[269] Ilan Hen. MIMO architecture for wireless communication. Intel Technology Journal, 10(2):157–165, May 15, 2006. ISSN 1535-

766X. URL http://developer.intel.com/technology/itj/2006/
volume10issue02/art07_MIMO_Architecture/p01_abstract.htm.

Chao:2006:P

[270] Lin Chao. Preface. Intel Technology Journal, 10(3):iii–iv, August 10, 2006.

ISSN 1535-766X.

Uhlig:2006:F

[271] Rich Uhlig. Forward: Intel Virtualization Technology: Taking virtualization mainstream on Intel architecture platforms. Intel Technology Journal, 10(3):v–vi, August 10, 2006. ISSN 1535-766X. URL http://developer.

intel.com/technology/itj/2006/v10i3/foreword.htm.

Anonymous:2006:TR

[272] Anonymous. Technical reviewers. Intel Technology Journal, 10(3):vii, August 10, 2006. ISSN 1535-766X.

Neiger:2006:IVT

[273] Gil Neiger, Amy Santoni, Felix Leung, Dion Rodgers, and Rich Uhlig. Intel Virtualization Technology: Hardware support for efficient processor virtualization. Intel Technology Journal, 10(3):167–177, August 10, 2006. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2006/v10i3/1-hardware/1-abstract.htm.

Abramson:2006:IVT

[274] Darren Abramson, Jeff Jackson, Sridhar Muthrasanallur, Gil Neiger, Greg Regnier, Rajesh Sankaran, Ioannis Schoinas, Rich Uhlig, Balaji Vembu, and John Weigert. Intel Virtualization Technology for directed I/O. Intel Technology Journal, 10(3):179–192, August 10, 2006.

ISSN 1535-766X. URL http://developer.intel.com/technology/
itj/2006/v10i3/2-io/1-abstract.htm.

Dong:2006:EXI

[275] Yaozu Dong, Shaofan Li, Asit Mallick, Jun Nakajim, Kun Tian, Xuefei Xu, Fred Yang, and Wilfred Yu. Extending Xen with Intel virtualization technology. Intel Technology Journal, 10(3):193–203, August 10, 2006. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2006/v10i3/3-xen/1-abstract.htm.

Ramachandran:2006:NCV

[276] Mahendra Ramachandran, Ned Smith, Matthew Wood, Sharad Garg, Jim Stanley, Eswar Eduri, Rinat Rappoport, Arie Chobotaro, Carl Klotz, and Lori Janz. New client virtualization usage models using Intel Virtualization Technology. Intel Technology Journal, 10(3):205–216, August 10, 2006. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2006/v10i3/4-models/1-abstract.htm.

Neumann:2006:IVT

[277] Dean Neumann, Dileep Kulkarni, Aaron Kunze, Gerald Rogers, and Edwin Verplanke. Intel Virtualization Technology in embedded and communications infrastructure applications. Intel Technology Journal, 10(3), August 10, 2006. ISSN 1535-766X. URL http://developer.intel.com/

technology/itj/2006/v10i3/5-communications/1-abstract.htm.

Fabian:2006:VE

[278] Patrick Fabian, Julia Palmer, Justin Richardson, Mic Bowman, Paul Brett, Rob Knauerhase, Jeff Sedayao, John Vicente, Cheng-Chee Koh, and Sanjay Rungta. Virtualization in the enterprise. Intel Technology Journal, 10(3):227–242, August 10, 2006. ISSN 1535-766X. URL http:/

/developer.intel.com/technology/itj/2006/v10i3/6-enterprise/
1-abstract.htm.

Casazza:2006:RSP

[279] Jeffrey P. Casazza, Michael Greenfield, and Kan Shi. Redefining server performance characterization for virtualization benchmarking. Intel Technology Journal, 10(3):243–251, August 10, 2006. ISSN 1535-766X.

REFERENCES 57

URL http://developer.intel.com/technology/itj/2006/v10i3/7-
benchmarking/1-abstract.htm.